### v1

# Improving Parallel Program Performance Through DSL-Driven Code Generation with LLM Optimizers

Abstract
Mapping computations to processors and assigning data to memory are critical for maximizing performance in parallel programming. These mapping decisions are managed through the development of specialized low-level system code, called mappers, crafted by performance engineers. Each mapper is tailored to a specific application and optimized for the underlying machine architecture, a process that requires days of refinement and tuning from an expert. Despite advances in system research, automating mapper generation remains a challenge due to the complexity of making millions of decisions to find the optimal solution and generate the solution as code. We introduce an approach that leverages recent advances in LLM-based optimizers for mapper design. In under ten minutes, our method automatically discovers mappers that surpass human expert designs in scientific applications by up to 1.34×1.34$\times$1.34 × speedup. For parallel matrix multiplication algorithms, our mapper achieves up to 1.31×1.31$\times$1.31 × of the expert-designed solution. To achieve this, we simplify the complexity of low-level code generation by introducing a domain-specific language (DSL) that abstracts the low-level system programming details and defines a structured search space for LLMs to explore. To maximize the application performance, we use an LLM optimizer to improve an agentic system that generates the mapper code. As a result, this approach significantly reduces the workload for performance engineers while achieving substantial performance gains across diverse applications. Finally, our results demonstrate the effectiveness of LLM-based optimization in system design and suggest its potential for addressing other complex system challenges.
1 Introduction
Task-based programming (Slaughter et al., 2015; Bauer et al., 2012; Augonnet et al., 2009; Chamberlain et al., 2007; Kaiser et al., 2014; Heller et al., 2017; Chandra et al., 2001; Duran et al., 2011; Moritz et al., 2018; Barham et al., 2022) has recently emerged as a prominent paradigm in parallel programming. The core idea is to decompose computations into self-contained functions called tasks that do not communicate with other tasks except through their arguments (typically tensors or multi-dimensional arrays). A performance-critical aspect of executing task-based applications is mapping, which involves assigning tasks to processors, mapping data to physical memories, and managing other low-level hardware resources. A concrete mapping policy, implemented as code, is referred to as a mapper. Mappers focus solely on improving performance without altering the correctness of the application’s output. The performance impact of a good versus a poor mapper can be substantial, often resulting in differences of an order of magnitude (Abdel-Gawad et al., 2014; Galvez et al., 2017).
Currently, writing mappers is a manual, labor-intensive, and time-consuming process that requires deep knowledge of the application, the machine, and the low-level system’s C++ mapping APIs. Even for experienced engineers, developing a high-performance mapper can take several days, requiring multiple rounds of refinement (SFX Teixeira et al., 2023).
In this paper, we focus on using LLMs to automatically generate mappers. There are two main challenges in using LLMs for mapper code generation. First, given a specific mapping policy described in natural language, LLMs struggle to generate the corresponding mapper code, which consists of a few hundred lines of low-level C++ code. This code generation task is significantly different from function-level code generation, which is more isolated and modular, where LLMs have demonstrated strong performance (Chen et al., 2021; Austin et al., 2021). In our case, generating C++ mappers requires a deep understanding of low-level systems and broader contextual dependencies in large-scale system code, which is beyond the capabilities of current LLMs (Jimenez et al., 2023; Du et al., 2023; Wang et al., 2024). The second challenge lies in identifying the optimal mapping strategy to maximize application performance. Many design problems in systems are naturally constrained optimization problems. Generating mappers that are syntactically correct is insufficient.
To address the first challenge, we design a Domain-Specific Language (DSL) for mapper generation. The DSL provides a high-level interface that encapsulates all the critical decisions a mapper has to make without interacting with the intricate low-level C++ APIs. The DSL defines a search space in its abstractions, which breaks a monolithic code generation down to generating multiple independent, modular components, which is much easier for code generation due to less contextual dependence between components.
For the second challenge, oirst, structuring the search space through the DSL’s decomposition and modularization of the mapper, and second, applying reinforcement learnformulating the search problem as an optimization problem where the objective is to maximize the performance of the parallel program. Similar approaches have achieved significant results in discovering faster algorithms for matrix multiplication (AlphaTensor) (Fawzi et al., 2022) and generating mathematical proofs (AlphaGeometry) (Trinh et al., 2024), both require searching through a discrete space for an optimal solution.
Recently, a great deal of work has demonstrated that instruction-tuned LLMs have an inherent ability to “optimize” for an objective, such as minimizing a function (Nie et al., 2024; Cheng et al., 2024), prompt optimization (Yang et al., ; Khattab et al., 2023), and molecule design optimization (Yuksekgonul et al., 2024). We propose a fully automated agent-based system to generate mappers and optimize for parallel application performance, illustrated in Figure 1. Our solution allows LLMs to refine their mapping strategies efficiently, reducing the time required to develop high-performance mappers from days to just a few minutes.
Figure 1: We illustrate an agent-based optimization for high-performance mapper code generation, where the mapper is written in a Domain-Specific Language (DSL) and iteratively refined by an agent through performance feedback. Initially, all tasks are mapped to the CPU and system memory. The LLM then improves by mapping computation-intensive tasks to the GPU and GPU’s FrameBuffer memory to minimize task execution time. Finally, tasks that require significant data exchange between the CPU and GPU are mapped to the ZeroCopy memory, enabling shared access and minimizing data movement overhead.
Our experiments show that mappers generated by LLMs not only match but can surpass expert-written mappers, achieving up to 1.34×1.34$\times$1.34 × speedup across 9 benchmarks. Notably, the search process enhances the performance of parallel matrix multiplication algorithms by up to 1.31×1.31$\times$1.31 ×. We also demonstrate, through ablation studies, the critical role of high-quality feedback in guiding LLM optimizers to discover more efficient mapping strategies. Our contributions are as follows:
  1. 1.
    Development of a Domain-Specific Language (DSL) to Simplify Mapper Generation: We design a DSL to address the complexities of direct code generation in low-level systems. This DSL provides a higher-level abstraction, enabling LLMs to generate mapping code more effectively by encapsulating the complexities of low-level C++ APIs.
  2. 2.
    Formulation of Mapper Generation as a Discrete Optimization Problem: We formulate the task of mapper code generation as a discrete optimization problem with the objective of maximizing application performance. Our DSL defines a more structured and constrained search space for LLM to explore.
  3. 3.
    Experimental Validation and Performance Comparison: Our DSL significantly improves the success rate of mapper code generation for a given mapping strategy from 0% in C++ to 80% in our DSL. Leveraging high-quality feedback, our agent-based solution discovers mappers that achieve up to 1.34×1.34$\times$1.34 × speedup compared with expert-designed mappers, enhancing the performance of parallel matrix multiplication algorithms by up to 1.31×1.31$\times$1.31 ×.
2 Related Work
Mapping in Parallel Programming Many parallel programming systems allow users to make their own mapping decisions, such as Legion (Bauer et al., 2012), StarPU (Augonnet et al., 2009; 2010), Chapel (Chamberlain et al., 2007), HPX (Kaiser et al., 2014; Heller et al., 2017), Sequoia (Fatahalian et al., 2006), Ray (Moritz et al., 2018), and Pathways (Barham et al., 2022). Several techniques have been proposed to automate mapping, including machine learning models (O’Boyle et al., 2013; Wang & O’Boyle, 2009), static analysis (Poesia et al., 2017; Ren et al., 2008), and auto-tuning (SFX Teixeira et al., 2023). Unlike previous work, we use an agent-based approach with LLMs, exploring a larger search space for mappers than traditional methods.
Automated Code Generation With the rise of LLMs, various models (Chen et al., 2021; Nijkamp et al., 2022; Li et al., 2023; Ouyang et al., 2022; Wei et al., 2023) have been trained on code, enabling them to assist with programming tasks and automate software development (Shypula et al., 2023; Yang et al., 2024b; Jain et al., 2024). While LLMs have shown success in generating isolated function-level Python code (Chen et al., 2021; Austin et al., 2021), they face challenges when generating system-level C++ code in large repositories due to complex APIs and broader contextual dependencies (Du et al., 2023; Wang et al., 2024). Thus, adopting LLMs to generate code in real-world software remains difficult (Jimenez et al., 2023). Our work addresses this by developing a domain-specific language (DSL) that provides a high-level abstraction to help LLMs generate mapper code more effectively.
Automated System Design Optimization The use of machine learning for optimizing system design has garnered significant attention in recent years. Learned cost models have proven effective in optimizing deep learning operators on GPUs (Zheng et al., 2020a; b; 2022b). Methods in reinforcement learning and multi-armed bandit have been applied to tasks such as chip floorplanning (Mirhoseini et al., 2021) and program autotuning (Ansel et al., 2014), automatic vectorization of programs (Haj-Ali et al., 2020a), and optimizing compiler phase ordering (Haj-Ali et al., 2020b). Previous work primarily focuses on hyperparameter configurations, while our work aims to improve system performance by directly generating mapper code.
Optimization with LLMs
Recently, LLMs have been increasingly applied to optimization problems traditionally found in numerical domains (Boyd & Vandenberghe, 2004). AhmadiTeshnizi et al. (2023) addressed structured optimization challenges, such as mixed-integer linear programming (LP), while Yang et al. (2024a) applied LLMs to unstructured problems, treating them as black-box optimizations. (Nie et al., 2024) highlighted the role of feedback in efficiently finding global minima for black-box functions. Cheng et al. (2024) formally defined optimization problems solvable by LLMs and proposed an optimizer based on execution graphs. More recently, Yuksekgonul et al. (2024) introduced an LLM optimizer inspired by gradient descent. These advancements demonstrate the potential of LLMs to optimize beyond task completion, improving performance metrics across various domains.
3 Mapper Generation Task
Task Definition
The problem we address is the automated generation of high-performance mappers for the Legion parallel programming framework (Bauer et al., 2012). Mappers determine how tasks are assigned to processors and how data is placed in memory to optimize performance. A well-designed mapper can achieve up to 10$\times$ speedup compared to random mapping strategies. However, mappers are highly context-specific and must be carefully tailored to an application’s input and the machine’s architecture. Finding an optimal mapper is akin to solving a combinatorial optimization problem—a process that is time-consuming, labor-intensive, and traditionally performed by experienced performance engineers. The search space for discovering the best mapping strategy is vast, growing to $2^{14}$ even for the simplest scientific applications (as shown in prior work with a smaller search space than ours (SFX Teixeira et al., 2023)), making it challenging to efficiently explore all possible solutions. Moreover, even with a clear mapping strategy, writing the corresponding mapper requires experts to produce hundreds of lines of low-level C++ code, a task that can take several hours. As a result, the entire performance tuning process can take several days to complete.
Figure 2: Mappers decide the placement of each task in the task graph to processors, the placement of data to memory, and how the iteration space of data is partitioned and mapped to different processors.
Mapping Decisions
Now we elaborate on the decisions that a mapper has to make. The first critical decision is the processor selection for each task, determining whether a task is better suited for GPUs, CPUs, or OpenMP runtime. This choice depends on factors such as task size, GPU memory, and kernel launch latency. For example, tiny tasks that require very little computation may prefer to run on CPUs due to the GPU kernel launch overhead, whereas tasks with large memory footprints might prefer to be assigned to OpenMP or CPU when GPU memory is insufficient.
Next, the memory placement of data across different memory spaces is crucial to performance. A mapper must decide where to place data — in the GPU’s FrameBuffer for faster access, in ZeroCopy memory for shared access between CPU and GPU, or in CPU system memory for large data. Each choice introduces a trade-off between memory access speed, memory usage, and transfer overhead.
The memory layout option determines the optimal memory arrangement for data structures. Depending on the memory access patterns and the underlying hardware, selecting between a Struct of Arrays (SOA) or an Array of Structures (AOS) layout, along with constraints on array ordering (Fortran-order or C-order) and memory alignment, can significantly impact performance due to cache efficiency and data locality. Such choice is task-dependent and processor-dependent.
Finally, mappers need to decide how to perform index mapping. As shown in Figure 2, index mapping controls how the data (or more accurately, parallel for loops of task launches) is partitioned and mapped to the processors. The data (or tensors) can be multi-dimensional, and the distributed machines can also be viewed as a processor space (nodes and processors within the node). Index mapping controls the mapping between the two index spaces. Prior work (Unger et al., 2022; Zheng et al., 2022a) shows that searching over how data is partitioned and mapped can significantly change inter-processor communication volume, which affects performance. Commonly-used index mapping functions are shown in Section A.3.
4 A New Approach: Generation and Optimization in a DSL
4.1 Design a Domain-Specific Language
We designed a domain-specific language for mapper generation. We distill the key performance-critical aspects of mapping into language constructs that are easily expressed in the DSL. By providing a higher-level abstraction than C++ APIs, the DSL simplifies interfacing with LLMs. Also, such a modular design enables LLMs to efficiently address the complex optimization challenges of generating high-performance mappers. To implement our DSL, we develop a compiler that can translate the mapper written in DSL into low-level C++ mapping APIs.
Figure 3(a) presents an example DSL mapper that illustrates the key features of our domain-specific mapping language. By contrast, Figure 3(b) shows the (already simplified) code snippet extracted from a C++ mapper. Notably, the C++ snippet covers only part of the functionality that the IndexMap statement (which we discuss in detail below) in the DSL provides.
While a typical C++ mapper requires approximately 400 lines of code, the corresponding DSL mapper can achieve the same functionality in just about 30 lines on average — reflecting a 14×14$\times$14 × reduction, as detailed in Table 1 (based on the 9 applications used in Section 5.2 and Section 5.3). This substantial reduction in code complexity makes the DSL a more accessible target for LLM-based code generation, as it abstracts away many low-level implementation details.
1# Map task0 to GPU.
2Task task0 GPU;
3
4# Place certain data onto GPU ZeroCopy.
5Region * ghost_region GPU ZCMEM
6
7# Specify layout in memory
8# (aligned to 64 bytes)
9Layout * * * C_order SOA Align==64
10
11# Define a cyclic mapping strategy
12def cyclic(Task task):
13 ip = task.ipoint;
14 mgpu = Machine(GPU);
15 node_idx = ip[0] % mgpu.size[0];
16 gpu_idx = ip[0] % mgpu.size[1];
17 return mgpu[node_idx, gpu_idx];
18
19IndexTaskMap task4 cyclic
(a) An example mapper in DSL
1void slice_task(const Task& task,
2 const SliceTaskInput &input,
3 SliceTaskOutput &output) {
4 vector<Processor> targets =
5 this->select_targets_for_task(ctx, task);
6 DomainT<2> space = input.domain;
7 Point<2> num_points =
8 space.bounds.hi - space.bounds.lo + ones;
9 Rect<2> blocks(zeroes, num_blocks - ones);
10 ...// **126 lines of C++ code ommitted**
11 for (PointInRectIterator<2> it(blocks); it() != NULL; it++) {
12 DomainT<2,coord_t> slice_space;
13 TaskSlice slice;
14 slice.domain = {slice_lo, slice_hi};
15 slice.proc = targets[index++ % targets.size()];
16 output.slices.push_back(slice);
17 }
18}
(b) Simplified code snippet from a C++ mapper
Figure 3: A DSL mapper and a short code snippet from C++ mapper. The C++ mapper has already been simplified, and the shown C++ snippet only contributes to part of the functionality achieved by the IndexTaskMap statement on Line 19 in the DSL mapper.
Next, we explain the DSL’s design, highlighting the immense search space it opens for optimization, which makes generating high-performance mappers a challenging task. The grammar of the DSL is shown in Section A.1.
The Task statement (Line 2) performs processor selection for tasks. Although it may seem straightforward to always prefer GPUs, this decision is complicated by other factors such as limited GPU memory and kernel execution time as explained in Section 3. This is a per-task decision.
The Region statement (Line 5) performs memory placement for data. Possible choices include GPU’s FrameBuffer or ZeroCopy memory, CPU’s System memory or RDMA memory. Such decisions need to be made for each argument of each task, which opens up a huge trade-off space between memory usage, task execution time, and data transfer costs as explained in Section 3.
The Layout statement (Line 9) defines memory layouts, supporting both SOA (Struct of Arrays) and AOS (Array of Structures), as well as array ordering constraints like C_order, F_order, and memory alignment. Optimizing memory layout is crucial for performance, as access patterns vary with tasks, processors, and data structures. This is a per-task, per-data, per-processor decision.
The IndexTaskMap statement (Line 19) enables the mapping of tasks to processors using custom functions, such as cyclic or block mapping strategies. This creates a mapping between the index space defined in application code (e.g., for loops) and the processor space of distributed machines. Many parallel operators (e.g., matrix multiplications) are based on such loops, and deciding how to map them to distributed machines is inherently difficult. The DSL enables users to express arbitrary arithmetic mappings between two index spaces, significantly increasing the complexity of the search space. Additionally, we introduce primitives that transform the processor space, allowing for more flexible and precise index mapping decisions, as detailed in Section A.2.
Application 1 2 3 4 5 6 7 8 9 Avg.
LoC in C++ 347 306 379 447 437 430 428 433 448 406
LoC in DSL 16 14 16 38 38 38 33 38 32 29
LoC Reduction $\times2$ $\times2$ 24×24$\times$24 × 12×12$\times$12 × 12×12$\times$12 × $\times1$ 13×13$\times$13 × $\times1$ 14×14$\times$14 × 14×14$\times$14 ×
Table 1: Lines of Code (LoC) of mapper written in DSL, and LoC reduction compared with C++
4.2 Learning to Generate via Interactive Feedback
Figure 4: We show the high-level design of a mapper agent. The agent takes two inputs: server-related specifications and application-related information. When a mapper is generated by the agent, we run the mapper code on the server and get performance feedback. Then, we use an LLM to update each code block to improve the performance.
1class MapperAgent(trace.Module):
2
3 @trace.bundle(trainable=True)
4 def gen_task_stmt(self, app) -> str:
5 """
6 Generate the policy for placing tasks.
7 Example generated code:
8 Task * GPU; Task task1 CPU;
9 """
10 code = ""
11 for task in app.tasks:
12 proc = random.choice(["GPU","CPU"])
[…]
14 return code
15
16 def forward(self, app):
17 task_stmt = self.gen_task_stmt(app)
18 region_stmt = self.gen_region_stmt(app)
19 layout_stmt = self.gen_layout_stmt(app)
20 ...
21 return task_stmt + region_stmt + ...
(a) The decision procedures in generating a mapper.
1policy = MapperAgent()
2params = policy.parameters()
3optimizer = trace.Optimizer(params)
4
5app = GetApplicationInfo()
6test = GetMapperEvaluator(app)
7
8for i in range(iterations):
9 # Forward pass
10 try:
11 mapper = policy(app)
12 # feedback (str) contains performance
13 feedback = test(mapper)
14 except TraceExecutionError as e:
15 feedback = str(e)
16 target = e.exception_node
17
18 # Backward pass and update
19 optimizer.zero_feedback()
20 optimizer.backward(target, feedback)
21 optimizer.step()
(b) Learning policy using LLM-based optimizers.
Figure 5: Example agent built with the interface offered by Trace. Figure 5(a) shows the mapper agent (each decision is defined as a function). Figure 5(b) shows how we incorporate the feedback from the execution to update the agent. The full agent code is shown in Figure A6.
We conceptualize the mapper generation problem as an online optimization problem. Given a triplet $(\Theta,\omega,{\mathcal{T}})$, where $\Theta$ is the overall set of possible code that can be produced to generate a mapper; $\omega$ defines the optimization objective, which is to create a mapper that maximizes the throughput; and ${\mathcal{T}}:\Theta\rightarrow{\mathcal{F}},{\mathcal{G}}$ is the task that returns $f,g={\mathcal{T}}(\theta)$, where $f$ is the feedback from executing the generated mapper (the performance metric measured after using the generated mapper to serve the application code), and $g$ is the process graph of how the mapper is generated. In our particular setup, the performance of the mapper (throughput) is deterministic because system researchers have carefully controlled all possible randomness in the environment. If the parameter space is numerical, this online optimization problem can be solved through bandit algorithms (Lattimore & Szepesvári, 2020), reinforcement learning (Sutton & Barto, 2018), or Bayesian optimization (Snoek et al., 2012), but these methods are less effective when the parameter is text (mapper code).
In this online optimization problem, we leverage the DSL to bring structure to the parameter space and make the online optimization more efficient. $\theta$ is the program code, $\omega$, and $f$ are expressed as text. We leverage an LLM as an optimizer to solve this optimization problem. LLM has been shown to have some ability to “optimize” by iteratively proposing better solutions when an explicit objective is described in the text. This phenomenon has been observed and leveraged by Yang et al. (2024a); Cheng et al. (2024); Yuksekgonul et al. (2024); Patel et al. (2024) to solve various optimization problems. We design our mapper agent to make various mapping decisions, from selecting processors for tasks to mapping index tasks onto the processor space. A high-level schema of the mapper agent is shown in Figure 4.
Agent Design
We express our mapper agent as a Python program (MapperAgent in Figure 5(a)). This program serves as the initial starting point of the optimization. In this mapper agent, we manually break down the task of generating a monolithic mapper into several independent code segments, allowing the agent to decide what code to generate for each segment. This strategy works because the design of our DSL removes unnecessary dependence between code segments. Our modularization approach mimics the least-to-most prompting (Zhou et al., 2022), and our efforts to use programs as prompts echo the program-as-thought approach (Chen et al., 2022).
Optimization
We illustrate our workflow design in Figure 5, with the full template in Figure A6. Using the framework Trace (Cheng et al., 2024), we only ask the LLM to generate the method decorated with @trace.bundle(trainable=True) at each iteration of optimization. During an optimization step (shown in Figure 5(b)), the LLM receives complete information about how the agent generated the previous mapper, as well as feedback from the system execution. The LLM optimizer needs feedback from the system to perform the optimization. The system feedback can be categorized into three types: (1) Compile Error, indicating a failure of the LLM to generate code capable of producing a valid DSL; (2) Execution Error, such as an application running out of GPU memory; and (3) Performance Metric, which refers to the application’s running time or throughput. The goal of the optimization process is to minimize running time, thereby increasing throughput. Examples are provided in the “System Feedback” column in Table 2.
Mapper System Feedback Enhanced Feedback
Explain Suggest
mapper1 Compile Error: Syntax error, unexpected :, expecting { N/A There should be no colon : in function definition.
mapper2 Execution Error: Assertion failed: stride does not match expected value. Memory layout is unexpected. Adjust the layout constraints or move tasks to different processor types.
mapper3 Performance Metric: Execution time is 0.03s. N/A Move more tasks to GPU to reduce execution time.
Table 2: System feedback and enhanced feedback for different mappers.
Enhanced Feedback
To accelerate the LLM’s search for high-performance mappers, we introduce enhanced feedback in addition to system feedback. Table 2 provides examples with enhanced feedback, with further examples available in Section A.7. There are two types of enhanced feedback: the first is to explain execution errors, providing more informative failure messages to guide LLMs in generating correct mapper; the second is to suggest modifications to mapping decisions to avoid errors or improve performance. Enhanced feedback is implemented via keyword matching, where system feedback triggers the corresponding explanations and suggestions for mapper improvements. We conduct an ablation study in Section 5.4 to demonstrate the effectiveness of enhanced feedback in our experiments.
5 Evaluation
Experiments are conducted on a GPU cluster where each node has two Intel 10-core E5-2640 v4 CPUs, 256G main memory, and four NVIDIA Tesla P100 GPUs. Regarding the LLM, we use gpt-4o-2024-08-06.
5.1 Effectiveness of the DSL
To evaluate the necessity and effectiveness of our DSL in mapper code generation, we compared the generation of mappers in C++ with DSL. We devised 10 mapping strategies, each described in natural language, to serve as test cases for code generation [^1] An example strategy is “aligning all data to 64 bytes in memory and utilizing Fortran ordering for multi-dimensional data.”. A complete list of these strategies is provided in Section A.9. It is important to note that the objective is not to optimize the performance of any specific application, but rather to assess whether the LLM can accurately generate a C++/DSL mapper based on a given strategy.
We provided the same types of materials in the prompt to ensure a fair comparison. We provided documentation [^2] We used documentation covering all features of DSL, and pre-existing documentation for C++ mapper., example mapper programs, and starting code for both DSL and C++ interfaces. We measured the success rate of generating mappers that can pass compilation and pass test cases for each strategy. For C++ code generation, we enhanced the process by incorporating compiler feedback for additional attempts. The LLM received compiler error messages and iteratively refined the code up to 10 times. We use DSPy (Khattab et al., 2023) to build our generation pipeline.
Code Generation Target Mapping Strategy Success Rate
1 2 3 4 5 6 7 8 9 10
C++ (single trial) 0%
C++ (iterative refine) 0%
DSL (single trial) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 80%
Table 3: Compilation and Strategy Test. We show the success rate for code generation given 10 mapping strategies. Generating DSL code significantly outperforms generating C++ (either without or with compiler feedback). – fails to compile, ✗ compiles but fails the test, and $\checkmark$ passes the test.
ResultTable 3 presents the performance of the LLM in mapper code generation in three settings: C++ single trial (without compiler feedback), C++ with compiler feedback, and DSL. We evaluate the LLM with 10 different mapping strategies. As observed, the DSL approach consistently outperforms the other settings, followed by C++ with compiler feedback, while C++ without feedback performs the worst. These results confirm that the LLM is poor at generating system-level C++ code and highlight the effectiveness and necessity of our DSL design in significantly enhancing the code generation capabilities for mapping.
Failure Case Analysis The compilation errors in C++ mapper generation arise because LLMs are unable to manage the framework-specific contextual dependencies inherent to low-level system software programming. For instance, LLMs generate variable names or references that do not exist within the provided codebase. These issues are particularly challenging to resolve since we cannot supply the full header files to the LLMs due to token limits.
In the cases where the code compiles but fails the test cases, the failures are exclusively in C++. The root cause lies in the complexity of implementing a C++ mapper, which requires a deep understanding and reasoning about the low-level mapping APIs. As outlined in the documentation, multiple APIs must work in concert to accomplish specific tasks, yet LLMs are unable to grasp this level of coordination. While compiler feedback can guide LLMs in avoiding trivial errors (e.g., such as generating non-existent variable names), it cannot bridge the gap in understanding the intricacies of low-level C++ mapping APIs. This limitation explains why all attempts at C++ code generation ultimately fail.
In contrast, only two test cases fail in the DSL context on the single trial, both due to compilation errors stemming from incorrect usage of the DSL’s syntax. This is understandable given that the DSL is a newly designed language with limited training data available for the LLM. Thus, for performance optimization experiments, we do not attempt to have LLMs optimize over C++ mappers, as results sufficiently demonstrate that LLMs struggle to generate C++ code that meets the specifications of what a desired mapper should do, let alone explore a large search space of different mapping strategies.
Figure 6: Normalized throughput for scientific applications comparing expert-written mappers, random mappers, best mappers found by Trace, and the average optimization trajectories of Trace and OPRO in 10 iterations across 5 runs.
5.2 Accelerating Scientific Applications
We use the following three scientific applications as our benchmarks. The circuit simulation benchmark (Bauer et al., 2012) models the behavior of an electrical circuit by simulating currents and voltages across interconnected nodes and wires. The stencil computation benchmark (Van der Wijngaart & Mattson, 2014) simulates a 2D grid where each point’s value is updated based on its neighbors using a stencil pattern. The Pennant benchmark (Ferenbaugh, 2015) models unstructured mesh, Lagrangian staggered-grid hydrodynamics, used for simulating compressible flow.
In this experiment, we compared the performance of different mappers. All mappers are implemented using our DSL. Our goal is to determine whether LLMs can effectively explore the search space of mappers by generating high-performance DSL mappers.
Expert-written mappers were manually developed and optimized by domain experts as part of the application development process. We re-implemented these expert-written C++ mappers using our DSL to establish a ground truth for comparison with our approach. Validation confirmed that the DSL-based mappers achieve performance equivalent to the original C++ mappers, providing a fair basis for evaluating the performance of our generated mappers.
Randomly generated mappers are produced by our MapperAgent with 10 different random seeds and serve as a baseline for this experiment. We ran 10 randomly generated mappers and reported the average performance.
LLM-generated mappers are implemented using Trace (Cheng et al., 2024), which employs LLMs as optimizers. We tested both the Trace and OPRO (Yang et al., ) search algorithms, running 10 iterations for each application. Due to the stochastic nature of LLM outputs, we repeated the optimization process 5 times and averaged the results. We also report the best mapper found by Trace across the 5 runs.
The performance-critical mapping decisions for these applications are the processor type selection, memory placement of data, and the layout selection for data. Other decisions available to mappers do not change the performance of these applications much. The simplest application above (with the smallest search space) is Stencil, which contains 2 tasks and 12 data arguments. Each task and data argument has two choices of placement, together with 4 additional layout choices for each data argument, forming an optimization space of $2^{38}$.
Results
We use normalized throughput as the performance metric in Figure 6, where higher values indicate better performance. The throughput is normalized w.r.t the expert-written mappers. As shown in Figure 6, random mappers are consistently the least effective across all applications, highlighting the importance of mapping decisions on application performance. When comparing the optimization trajectories of Trace and OPRO, Trace performs similarly to OPRO, and is slightly better than OPRO on Pennant. All the best mappers found by Trace can at least match the performance of expert mappers.
Figure 7: Normalized throughput for matrix-multiplication algorithms. We compare expert-written mappers, random mappers, best mappers found by Trace, and the average optimization trajectories of Trace and OPRO in 10 iterations across 5 runs.
Interestingly, the best mapper identified in the Circuit benchmark outperforms the expert mapper by 1.34×1.34$\times$1.34 ×. Upon manual investigation, we observed that the key difference lies in memory placement: the best mapper allocates two data collections to the GPU FrameBuffer memory, whereas the expert mapper places these collections in GPU ZeroCopy memory. This strategy reduces task execution time, despite a slight increase in inter-GPU communication costs, ultimately leading to improved overall performance. For the Pennant benchmark, while there is a minor difference in data collection placement, the final performance results between the two mappers are nearly equivalent.
For each application, the optimization process completes within 10 minutes, significantly reducing mapper development time from days to minutes. This substantial improvement highlights the efficiency of our LLM-enhanced DSL in quickly generating high-performance mappers, offering clear benefits for both developers and application performance.
5.3 Accelerating Matrix Multiplications
We follow the same experimental setup as described in Section 5.2, with the key difference being the focus on matrix multiplication algorithms in this subsection. Different from the applications in Section 5.2, the important mapping decision here is index mapping, which is the choice of how to launch concurrent tasks and distribute the tiles of matrices across multiple GPUs.
In the DSL mapper, achieving a high-performance index mapping requires specifying a function that maps the iteration space (defined by the parallel for loop) to the processor space of the machine (discussed in Section 4.1 IndexTaskMap). Based on our estimates, each algorithm has approximately $10^{9}$ possible choices for index mapping. We show some of the mapping functions in Section A.5
We target 6 different matrix multiplication algorithms: Cannon’s (Cannon, 1969), SUMMA (Van De Geijn & Watts, 1997), PUMMA (Choi et al., 1994), Johnson’s (Agarwal et al., 1995), Solomonik’s (Solomonik & Demmel, 2011), and COSMA (Kwasniewski et al., 2019). Each algorithm exhibits different performance characteristics and may be the preferred implementation of matrix multiply depending on the target machine and input size. Our goal is to explore and identify better mappings for these algorithms. We elaborate further on these algorithms in Section A.4.
Results
As shown in Figure 7, the algorithm self-specified expert mappers’ compute throughput is normalized to 1.0. Random mappers only yield 2%-40% of the throughput achieved by expert mappers, highlighting the importance of high-quality mapper design. The best mappers found by Trace consistently outperform the expert mappers, achieving speedups ranging from 1.09×1.09$\times$1.09 × to 1.31×1.31$\times$1.31 ×. For both PUMMA and Solomonik’s, the throughput of the best mapper discovered by Trace is significantly higher than the average optimization trajectory across 5 runs. This variability is due to the randomness of LLMs in our experiments, where even small changes in mappers can lead to notable performance differences. When comparing optimization trajectories of Trace and OPRO, Trace shows significant performance gains in SUMMA, PUMMA, Johnson’s, and COSMA, while achieving similar results on other benchmarks.
Finally, we inspected and compared the mapping decisions of the expert mappers with those of the best mappers found through our search. The performance improvements are entirely attributed to more effective index mapping. Index mapping controls the partitioning and distribution of data, in this case, matrices, across the GPUs. The optimized mapping reduces inter-GPU communication and enhances data locality, leading to improved performance in parallel matrix multiplication algorithms.
Figure 8: Comparison of three types of feedback design. System provides only the system execution information (performance numbers or error messages) to the optimizer. Explain provides additional explanations of execution errors. Suggest provides modification proposals to the mappers. All feedback is automatically generated.
5.4 Ablation Study of Feedback
Within the Trace framework, feedback is provided to the LLM optimizer to inform subsequent improvement on mappers. The quality of the feedback directly affects the success of the optimization process. We evaluate three types of feedback to assess how different feedback influences the performance of LLM-based optimizers during mapper search (described in Table 2): (1) system feedback (compile error, execution error, or performance metrics), (2) error explanations, and (3) suggestions for mapper adjustments.
The first feedback message includes only system feedback (labeled System in Figure 8). Next, we evaluate an enhanced feedback message that includes both system feedback and error explanations (labeled System+Explain). Finally, we provide the full feedback message, including all three types (labeled System+Explain+Suggest), corresponding to the Trace results shown in Figure 6 and Figure 7. We evaluate on three benchmarks: the Circuit simulation benchmark and two matrix multiplication algorithms (COSMA and Cannon’s).
As shown in Figure 8, across all three benchmarks, the full feedback message consistently achieves the highest throughput after 10 iterations, followed by the one without guidance. The system-only feedback performs the worst among the three. While the degree of impact from the feedback types varies across benchmarks, the results clearly demonstrate the critical role of high-quality feedback in guiding LLM optimizers to discover more efficient mappers.
6 Conclusion
In this paper, we addressed the challenges of automating mapper generation in task-based programming through the use of LLMs and a Domain-Specific Language (DSL). By designing a high-level DSL, we effectively simplified the complex task of generating low-level C++ mappers, enabling LLMs to handle mapper generation without deep knowledge of system intricacies. Additionally, we formulated the mapper generation task as a discrete optimization problem, leveraging LLM optimizers to explore a structured and constrained search space defined by the DSL. Our experimental results demonstrate the efficacy of this approach, with LLM-generated mappers achieving up to 1.34×1.34$\times$1.34 × speedup over expert-written mappers across nine benchmarks. For matrix multiplication algorithms, we observed a performance boost of up to 1.31×1.31$\times$1.31 ×. These findings show that our LLM-enhanced DSL significantly reduces development time from days to minutes, benefiting both human developers and the performance of parallel applications.
Reproducibility Statement
We provide comprehensive details regarding the setup of our benchmark, ensuring full reproducibility based on the provided information. We are planning to open-source the framework, as well as the code to reproduce all experiment results. We provide detailed documentation of the evaluation process, along with all the prompts that we use. We use gpt-4o-2024-08-06 with a max token of 16383 and default temperature. All software used in our evaluation (e.g., Legion) is open source. We acknowledge the potential difference in performance measurement across different platforms.
Acknowledgments
We thank Yuhui Zhang, Chenyang Yang, Genghan Zhang, Qizheng Zhang, Mohammad R. Fadiheh, Ed Chen, Mert Yuksekgonul, Jiaxin Ge, Tingyu Zhu, Simon Guo and Anne Ouyang for the feedback. We thank Ching-An Cheng and Ken Liu for their discussions and feedback.
Appendix A Appendix
A.1 DSL Grammar
We show the DSL grammar in Figure A1.
[…]
$\mathsf{IndexTaskMap}\;\mathbf{TaskName}\;\mathbf{var}\mathrel{\mid}$
$\mathsf{SingleTaskMap}\;\mathbf{TaskName}\;\mathbf{var}\mathrel{\mid}$
$\mathtt{FuncDef}\mathrel{\mid}\mathtt{TaskMapModifier}$
[…]
$\mathsf{InstanceLimit}\;\mathbf{TaskName}\;\mathbf{int}\mathrel{\mid}\ldots$
[…]
$\mathtt{Expr}\;\mathsf{Op}\;\mathtt{Expr}\mathrel{\mid}(\mathtt{Expr})\mathrel{\mid}\mathtt{Expr}[\mathtt{Expr}]\mathrel{\mid}*\;\mathtt{Expr}$
$\mathtt{Expr}\;\mathtt{?}\;\mathtt{Expr}:\mathtt{Expr}$
Figure A1: Grammar of DSL. A DSL program is a list of statements, each of which controls one aspect of mapping.
A.2 Processor Space Transformation for Index Mapping
Transformation Semantics
$m^{\prime}=m.\mathtt{split}(i,d)$ $m^{\prime}[a_{0},\ldots,a_{n+1}]\coloneq m[b_{0},\ldots,b_{n-1},b_{n}]$
$b_{t}=\begin{cases}a_{t}\qquad&\quad t<i\\ a_{i}+a_{i+1}\cdot d\qquad&\quad t=i\\ a_{t+1}\qquad&\quad t>i\end{cases}$
$m^{\prime}=m.\mathtt{merge}(p,q)$ $m^{\prime}[a_{0},\ldots,a_{n-1}]\coloneq m[b_{0},\ldots,b_{n-1},b_{n}]$
$p<q$ $b_{t}=\begin{cases}a_{t}\qquad&\quad t<p\vee p<t<q\\ a_{p}\%m.size[p]\qquad&\quad t=p\\ a_{p}/m.size[p]\qquad&\quad t=q\\ a_{p-1}\qquad&\quad t>q\\ \end{cases}$
$m^{\prime}=m.\mathtt{swap}(p,q)$ $m^{\prime}[a_{0},\ldots,a_{n-1}]\coloneq m[b_{0},\ldots,b_{n-1}]$
$b_{t}=\begin{cases}a_{q}\qquad&\quad t=p\\ a_{p}\qquad&\quad t=q\\ a_{t}\qquad&\quad t\neq p\land t\neq q\\ \end{cases}$
$m^{\prime}=m.\mathtt{slice}(i,low,high)$ $m^{\prime}[a_{0},\ldots,a_{n-1}]\coloneq m[b_{0},\ldots,b_{n-1}]$
$0\leq low\leq high<m.size[i]$ $b_{t}=\begin{cases}a_{i}+low\qquad&\quad t=i\\ a_{t}\qquad&\quad t\neq i\\ \end{cases}$
Figure A2: Semantics of processor space transformations expressed as mappings from the indices of the transformed processor space to the indices of the original processor space.
A processor space $m$ can be initialized via m = Machine(GPU); as a 2D tuple where the first dimension represents the node index, and the second dimension represents the index of GPU within the node. To better express mapping between index spaces, we define a set of transformation primitives. A transformation primitive is a function of the processor space $m$ that returns a transformed processor space $m^{\prime}$, where $m$ and $m^{\prime}$ are related through the mapping shown in the right-hand side of Figure A2.
Our transformations are inspired by widely used operations for changing the dimensionality of arrays in libraries such as NumPy, but the application to mapping is quite different. We now explain each transformation in detail.
The split transformation takes two arguments: an integer $i$ that indicates the dimension to be split, and the splitting factor $d$. Suppose $m$ is a processor space of size $(8,8)$, then after executing $m^{\prime}=m.\mathtt{split}(0,2)$, $m^{\prime}$ will be a processor space of size $(2,4,8)$. An important property of split and all DSL transformations is that they are invertible. Thus, mappers can work with the transformed space $m^{\prime}$ but DSL can translate such uses back into the original processor space $m$ to identify which concrete processors to use. In this example, $m^{\prime}[j_{0},j_{1},j_{2}]=m[j_{0}+j_{1}\times 2,j_{2}]$.
The merge transformation takes two dimensions of the original processor to be fused as its input. Suppose $m^{\prime}$ is a processor space of size $(2,4,8)$. After applying $m^{\prime\prime}=m^{\prime}.\mathtt{merge}(0,1)$, $m^{\prime\prime}$ will be a processor space of size $(8,8)$. The processor indexed by $m^{\prime\prime}[j_{0},j_{1}]$ corresponds to $m^{\prime}[j_{0}\%2,j_{0}/2,j_{1}]$.
Transformations can be chained together. Suppose $m$ is a 2D processor space, and we start with $m^{\prime}=m.\mathtt{split}(0,d)$, followed by $m^{\prime\prime}=m^{\prime}.\mathtt{merge}(0,1)$. The final processor space $m^{\prime\prime}$ is a 2D processor space. Now we will derive the index transformation from $m^{\prime\prime}$ to $m$ by applying the transformation rules from merge transformation and split transformation: $m^{\prime\prime}[j_{0},j_{1}]=m^{\prime}[j_{0}\%d,j_{0}/d,j_{1}]=m[(j_{0}\%d)+(j_{0}/d)\times d,j_{1}]$. The expression $(j_{0}\%d)+(j_{0}/d)\times d$ can be simplified to $j_{0}$ because the division operator /// between two integers rounds to zero. Therefore, $m^{\prime\prime}[j_{0},j_{1}]=m[j_{0},j_{1}]$, showing that the split and merge transformation primitives are inverses of each other.
The swap transformation primitive takes two parameters indicating the dimensions to be swapped and returns a processor space where the two indices of the chosen dimensions are flipped. The swap transformation is often combined with merge: The merge primitive linearizes two dimensions into one, but there is a choice whether to use row-major or column-major iteration order in the linearization. Users can change the iteration order for the merge by swapping the two dimensions.
The slice transformation primitive takes three parameters, the dimension to slice, and the lower bound and upper bound of the dimension. The index mapping rule for the slice transformation is to add a constant shift in the chosen dimension. The slice transformation primitive can be useful if users want to map the iteration space to only part of the original processor space. If two iteration spaces can be executed concurrently in the program, users can map one iteration space to half of the processors and map the other iteration space to the other half of the processors. In this case, the slice transformation allows users to map an iteration space to the selected portion of the processor space.
A.3 Common Index Mapping Functions
We show the common index mapping functions in Figure A3.
Figure A3: Common transformations and index mapping functions. The shaded subarea of the iteration space will be mapped to the shaded processor in the processor space. The transformation code can transform the original $(2,2)$ processor space into the desired processor space. The result processor space will be used for mapping in the user-defined function.
A.4 Parallel Matrix Multiplication Algorithms
2D Algorithms
Cannon’s (Cannon, 1969) introduced a systolic communication pattern with tiled data partitioning for distributed matrix multiplication. PUMMA (Choi et al., 1994) and SUMMA (Van De Geijn & Watts, 1997) extended this approach by supporting non-square matrices and improving communication efficiency through pipelining. They are called 2D algorithms because they partition the matrices into 2D tiles and then map them onto the processor space.
Non-2D Algorithms
Johnson’s (Agarwal et al., 1995) introduced a 3D algorithm that partitions the input matrices into 3D tiles and uses additional memory per processor to reduce communication compared to 2D algorithms. Solomonik’s (Solomonik & Demmel, 2011) balances between 2D and 3D approaches by using extra memory to further minimize communication. COSMA (Kwasniewski et al., 2019) takes a different approach by optimizing the processor grid and parallelization strategy based on the input size and the machine size.
A.5 Index mapping functions used by Matrix Multiplication Algorithms
We show some index mapping functions used by matrix multiplication algorithms in Figure A4
Figure A4: Example mapping functions used by the mappers of matrix multiplication algorithms.
A.6 Explanation of Index Mapping for Solomonik’s Algorithm
Figure A5: Mapper of the Solomonik’s algorithm on a 2-node machine with 4 GPUs per node. The 2D processor space is transformed via the split transformation primitive into a 6D space (visualized as two 3D spaces).
Figure A5 shows a mapper for the Solomonik’s algorithm on a 2-node machine with 4 GPUs per node. The result distribution for the 3D iteration space is that each node will get half of the whole iteration space by partitioning along the x-axis, and the 4 GPUs per node will perform a 2D block distribution over the y-z plane.
There is a dimension mismatch between the iteration space (3D) and the initial processor space (2D). To conduct the desired mapping required by the algorithm, we first apply the split transformation primitive four times (colored as red in the code). We apply the first (resp. last) two split transformations to make the node dimension (resp. GPU dimension) align with the 3D iteration space. We visualize the result 6D processor space as two 3D spaces. The first 3D space (representing the node dimension) is of size $(2,1,1)$ and the second 3D space (representing the GPU dimension) is of size $(1,2,2)$.
A.7 Examples of Feedback Configurations
We give examples for the system feedback and enhanced feedback in Table A1. The enhanced feedback includes explanations of errors and suggestions for mapper modifications.
Mapper System Feedback Enhanced Feedback
Explain Suggest
mapper1 Compile Error: Syntax error, unexpected :, expecting { N/A There should be no colon : in function definition.
mapper2 Compile Error: IndexTaskMap’s function undefined N/A Define the IndexTaskMap function first before using it.
mapper3 Compile Error: mgpu not found N/A Include mgpu = Machine(GPU); in the generated code.
mapper4 Execution Error: Assertion failed: stride does not match expected value. Memory layout is unexpected. Adjust the layout constraints or move tasks to different processor types.
mapper5 Execution Error: DGEMM parameter number 8 had an illegal value Memory layout is unexpected. Adjust the layout constraint.
mapper6 Execution Error: Slice processor index out of bound IndexTaskMap statements cause error. Ensure that the first index of mgpu ends with % mgpu.size[0], and the second element ends with % mgpu.size[1].
mapper7 Execution Error: Assertion ‘event.exists()’ failed InstanceLimit statements cause error. Avoid generating InstanceLimit statements.
mapper8 Performance Metric: Execution time is 0.03s. N/A Move more tasks to GPU to reduce execution time.
mapper9 Performance Metric: Achieved throughput = 4877 GFLOPS N/A Try using different IndexTaskMap or SingleTaskMap statements to maximize throughput.
Table A1: System feedback and enhanced feedback (error explanations and adjustment suggestions) for different mappers.
A.8 Trace Agent Code
Trace (Cheng et al., 2024) uses Python decorators like @bundle to annotate Python programs. It allows us to design an LLM code generation agent as if we were writing a Python program ourselves. We first set up an end-to-end runnable Python program that can generate a valid mapper program by randomly making decisions over the search space. We show the high-level structure of our Trace Mapper in Figure A6. At each optimization step, Trace will execute DSLMapperGenerator and collect the corresponding execution flow to build up a graph. Then it will make a call to an LLM to perform an update to any function that is decorated with @bundle(trainable=True). The DSLMapperGenerator is structured in the same way as providing a search space specified by the DSL, where an LLM optimizer can make decisions along the pre-designed axes. We note that this type of design is only enabled by recent developments like Trace and is much more challenging to do using older LLM-based frameworks.
1import opto.trace as trace
2
3class MapperAgent(trace.Module):
4 @trace.bundle(trainable=True)
5 def task_decision(self, tasks):
6 ...
7
8 @trace.bundle(trainable=True)
9 def region_decision(self, regions):
10 ...
11
12 @trace.bundle(trainable=True)
13 def layout_decision(self):
14 ...
15
16 @trace.bundle(trainable=True)
17 def instance_limit_decision(self, tasks):
18 ...
19
20 @trace.bundle(trainable=True)
21 def index_task_map_decision(self, index_tasks):
22
23 @trace.bundle(trainable=True)
24 def single_task_map_decision(self, single_tasks):
25 ...
26
27 def generate_mapper(self):
28 """
29 Generate the final mapper code by combining all code statements.
30 """
31 task_statements = self.task_decision(self.tasks)
32 region_statements = self.region_decision(self.regions)
33 layout_statements = self.layout_decision()
34 instance_limit_statements = self.instance_limit_decision(self.tasks)
35 index_task_map_statements = self.index_task_map_decision(self.index_tasks, self.index_task_specification)
36 single_task_statements = self.single_task_map_decision(self.single_tasks)
37
38 code_statements = (
39 task_statements +
40 region_statements +
41 layout_statements +
42 instance_limit_statements +
43 index_task_map_statements +
44 single_task_statements
45 )
46 # Combine all code statements and function definitions into a single string
47 code_list = code_statements
[…]
49 return mapper_code
Figure A6: High-level structure of the Trace-based agent template, where functions annotated with @bundle(trainable=True) define the search space that the LLM optimizer updates during mapper generation. Note: This agent serves as a shared starting point for ALL tasks. For each task, we produce a mapper from this starting agent and then ask LLMs to “optimize” this agent (by changing functions that are trainable) to produce mappers that are optimal for the particular task.
A.9 Mapping Strategies
Strategy 1: Map the tasks of calculate_new_currents, distribute_charge, update_voltages onto GPUs in this way: linearize the 2D GPU processor space into 1D, then perform 1D block mapping from launch domain to the linearized 1D processor space.
1Task * GPU,CPU; # for any task, run on GPU if supported
2Region * *GPU FBMEM; # for any task, any region, if mapped onto GPU, use FBMEM as default
3Region * * CPU SYSMEM; # if mapped onto CPU, use SYSMEM as default
4
5Layout * * * SOA C_order;
6
7mcpu = Machine(CPU);
8mgpu = Machine(GPU);
9
10========== Above is fixed ==========
11def linearblock(Task task) {
12 return mgpu[task.ipoint[0] / mgpu.size[1], task.ipoint[0] % mgpu.size[1]];
13}
14
15IndexTaskMap calculate_new_currents,distribute_charge,update_voltages linearblock;
Strategy 2: Place ghost/shared regions (rp_shared and rp_ghost) onto GPU zero-copy memory
1Task * GPU,CPU; # for any task, run on GPU if supported
2
3Region * * GPU FBMEM; # for any task, any region, if mapped onto GPU, use FBMEM as default
4Region * * CPU SYSMEM; # if mapped onto CPU, use SYSMEM as default
5
6Layout * * * SOA C_order;
7
8mcpu = Machine(CPU);
9mgpu = Machine(GPU);
10
11========== Above is fixed ==========
12
13Region * rp_shared GPU ZCMEM;
14Region * rp_ghost GPU ZCMEM;
Strategy 3: Use Array Of Struct (AOS) data layout for all data instead of the default SOA
1Task * GPU,CPU; # for any task, run on GPU if supported
2
3Region * * GPU FBMEM; # for any task, any region, if mapped onto GPU, use FBMEM as default
4Region * * CPU SYSMEM; # if mapped onto CPU, use SYSMEM as default
5
6mcpu = Machine(CPU);
7mgpu = Machine(GPU);
8
9========== Above is fixed ==========
10
11Layout * * * AOS;
Strategy 4: Use Fortran ordering of data layout for all data instead of the default C order
1Task * GPU,CPU; # for any task, run on GPU if supported
2
3Region * * GPU FBMEM; # for any task, any region, if mapped onto GPU, use FBMEM as default
4Region * * CPU SYSMEM; # if mapped onto CPU, use SYSMEM as default
5
6mcpu = Machine(CPU);
7mgpu = Machine(GPU);
8
9========== Above is fixed ==========
10
11Layout * * * F_order;
Strategy 5: Align all the regions to 64 bytes while using the Fortran ordering of data
1Task * GPU,CPU; # for any task, run on GPU if supported
2
3Region * * GPU FBMEM; # for any task, any region, if mapped onto GPU, use FBMEM as default
4Region * * CPU SYSMEM; # if mapped onto CPU, use SYSMEM as default
5
6mcpu = Machine(CPU);
7mgpu = Machine(GPU);
8
9========== Above is fixed ==========
10
11Layout * * * Align==64 F_order;
Strategy 6 Place the task calculate_new_currents onto CPU
1Task * GPU,CPU; # for any task, run on GPU if supported
2
3Region * * GPU FBMEM; # for any task, any region, if mapped onto GPU, use FBMEM as default
4Region * * CPU SYSMEM; # if mapped onto CPU, use SYSMEM as default
5
6mcpu = Machine(CPU);
7
8mgpu = Machine(GPU);
9
10Layout * * * SOA C_order;
11
12========== Above is fixed ==========
13Task calculate_new_currents CPU;
Strategy 7: Collect all the memory used by task calculate_new_currents
1Task * GPU,CPU; # for any task, run on GPU if supported
2
3Region * * GPU FBMEM; # for any task, any region, if mapped onto GPU, use FBMEM as default
4Region * * CPU SYSMEM; # if mapped onto CPU, use SYSMEM as default
5
6mcpu = Machine(CPU);
7mgpu = Machine(GPU);
8
9Layout * * * SOA C_order;
10
11========== Above is fixed ==========
12CollectMemory calculate_new_currents *;
Strategy 8: Ensure that at most 4 tasks of calculate_new_currents can be run at the same time
1Task * GPU,CPU; # for any task, run on GPU if supported
2
3Region * * GPU FBMEM; # for any task, any region, if mapped onto GPU, use FBMEM as default
4Region * * CPU SYSMEM; # if mapped onto CPU, use SYSMEM as default
5
6mcpu = Machine(CPU);
7mgpu = Machine(GPU);
8
9Layout * * * SOA C_order;
10
11========== Above is fixed ==========
12InstanceLimit calculate_new_currents 4;
Strategy 9: Map the second region argument of task distribute_charge onto GPU’s Zero-Copy memory
1Task * GPU,CPU; # for any task, run on GPU if supported
2
3Region * * GPU FBMEM; # for any task, any region, if mapped onto GPU, use FBMEM as default
4Region * * CPU SYSMEM; # if mapped onto CPU, use SYSMEM as default
5
6mcpu = Machine(CPU);
7mgpu = Machine(GPU);
8
9Layout * * * SOA C_order;
10
11========== Above is fixed ==========
12Region distribute_charge 1 GPU ZCMEM;
Strategy 10: Map the tasks of calculate_new_currents,distribute_charge,update_voltages onto GPUs in a 1D cyclic manner: perform a cyclic distribution over both the node and processor dimensions.
1Task * GPU,CPU; # for any task, run on GPU if supported
2
3Region * * GPU FBMEM; # for any task, any region, if mapped onto GPU, use FBMEM as default
4Region * * CPU SYSMEM; # if mapped onto CPU, use SYSMEM as default
5
6mcpu = Machine(CPU);
7mgpu = Machine(GPU);
8
9Layout * * * SOA C_order;
10
11========== Above is fixed ==========
12def cyclic1d(Task task) {
13 ip = task.ipoint;
14 # cyclic over node, cyclic over gpu
15 return mgpu[ip[0] % mgpu.size[0], ip[0] / mgpu.size[0] % mgpu.size[1]];
16}
17
18IndexTaskMap calculate_new_currents,distribute_charge,update_voltages cyclic1d;
A.10 Generated Mapper Examples
Here we provide examples of generated mappers for a subset of problems. The mappers, written in DSL, are produced by the mapper agent. While the LLM is responsible for creating and refining the mapper agent, the agent itself is implemented in Python, and it generates mappers as DSL programs. For the Circuit Simulation benchmark, the optimized mapper (Figure A8) is more concise than the initial version (Figure A7), with an additional constraint for byte alignment in the data layout. In contrast, for Solomonik’s algorithm, the initial mapper is relatively simple (Figure A9), whereas the final optimized mapper adopts a more complex and detailed index mapping strategy (Figure A10).
1Task * GPU,OMP,CPU;
2Task calculate_new_currents GPU;
3Task update_voltages GPU;
4Region * * GPU FBMEM;
5Region * * * SOCKMEM,SYSMEM;
6Region * all_times GPU FBMEM;
7Region * all_nodes GPU FBMEM;
8Region * all_wires GPU FBMEM;
9Region * ghost_ranges GPU FBMEM;
10Region * rp_all_nodes GPU FBMEM;
11Region * all_private GPU FBMEM;
12Region * all_shared GPU FBMEM;
13Region * rp_shared GPU FBMEM;
14Region * rp_wires GPU FBMEM;
15Region * rp_ghost_ranges GPU FBMEM;
16Layout * * * C_order AOS;
17mgpu = Machine(GPU);
18
19m_2d = Machine(GPU);
20def same_point(Task task) {
21 return m_2d[*task.parent.processor(m_2d)];
22}
Figure A7: For the Circuit task, we show the mapper produced by the mapper agent at iteration 2.
1Task * GPU,OMP,CPU;
2Task calculate_new_currents GPU;
3Task update_voltages GPU;
4Region * * GPU FBMEM;
5Layout * * * C_order AOS Align==128;
6mgpu = Machine(GPU);
7
8m_2d = Machine(GPU);
9def same_point(Task task) {
10 return m_2d[*task.parent.processor(m_2d)];
11}
Figure A8: For the Circuit task, we show the mapper produced by the mapper agent at iteration 10.
1Task * GPU,OMP,CPU;
2Region * * GPU FBMEM;
3Region * * * SOCKMEM,SYSMEM;
4Layout * * * F_order SOA;
5mgpu = Machine(GPU);
6
7def block1d(Task task) {
8 ip = task.ipoint;
9 return mgpu[ip[0] % mgpu.size[0], ip[0] % mgpu.size[1]];
10}
11
12IndexTaskMap task_2 block1d;
13
14m_2d = Machine(GPU);
15def same_point(Task task) {
16 return m_2d[*task.parent.processor(m_2d)];
17}
Figure A9: For Solomonik’s algorithm, we show the mapper produced by the mapper agent at iteration 2.
1Task * GPU,OMP,CPU;
2Region * * GPU FBMEM;
3Region * * * SOCKMEM,SYSMEM;
4Layout * * * C_order SOA No_Align;
5mgpu = Machine(GPU);
6
7def block1d(Task task) {
8 ip = task.ipoint;
9 return mgpu[ip[0] % mgpu.size[0], ip[0] % mgpu.size[1]];
10}
11
12IndexTaskMap task_1 block1d;
13
14def cyclic1d(Task task) {
15 ip = task.ipoint;
16 linearize = ip[0] * 2 + ip[1];
17 return mgpu[ip[0] % mgpu.size[0], linearize % mgpu.size[1]];
18}
19
20IndexTaskMap task_1 cyclic1d;
21
22def cyclic2d(Task task) {
23 ip = task.ipoint;
24 linearize = ip[0] + ip[1] * 2;
25 return mgpu[ip[0] % mgpu.size[0], linearize % mgpu.size[1]];
26}
27
28IndexTaskMap task_1 cyclic2d;
29
30def linearize3D(Task task) {
31 ip = task.ipoint;
32 linearize = ip[0] + ip[1] + ip[2];
33 return mgpu[linearize % mgpu.size[0], linearize % mgpu.size[1]];
34}
35
36IndexTaskMap task_1 linearize3D;
37
38def linearize2D(Task task) {
39 ip = task.ipoint;
40 linearize = ip[0] * 2 + ip[2];
41 return mgpu[linearize % mgpu.size[0], linearize % mgpu.size[1]];
42}
43
44IndexTaskMap task_1 linearize2D;
45IndexTaskMap task_2 block1d;
46IndexTaskMap task_2 cyclic1d;
47IndexTaskMap task_2 cyclic2d;
48IndexTaskMap task_2 linearize3D;
49IndexTaskMap task_2 linearize2D;
50IndexTaskMap task_3 block1d;
51IndexTaskMap task_3 cyclic1d;
52IndexTaskMap task_3 cyclic2d;
53IndexTaskMap task_3 linearize3D;
54IndexTaskMap task_3 linearize2D;
55IndexTaskMap task_5 block1d;
56IndexTaskMap task_5 cyclic1d;
57IndexTaskMap task_5 cyclic2d;
58IndexTaskMap task_5 linearize3D;
59IndexTaskMap task_5 linearize2D;
60
61m_2d = Machine(GPU);
62def same_point(Task task) {
63 return m_2d[*task.parent.processor(m_2d)];
64}
Figure A10: For Solomonik’s algorithm, we show the mapper produced by the mapper agent at iteration 10.

### v2

@@
-Improving Parallel Program Performance Through DSL-Driven Code Generation with LLM Optimizers
+Improving Parallel Program Performance with LLM Optimizers via Agent-System Interface
 Abstract
-Mapping computations to processors and assigning data to memory are critical for maximizing performance in parallel programming.
-These mapping decisions are managed through the development of specialized low-level system code, called mappers, crafted by performance engineers.
-Each mapper is tailored to a specific application and optimized for the underlying machine architecture, a process that requires days of refinement and tuning from an expert.
-Despite advances in system research, automating mapper generation remains a challenge due to the complexity of making millions of decisions to find the optimal solution and generate the solution as code.
-We introduce an approach that leverages recent advances in LLM-based optimizers for mapper design.
-In under ten minutes, our method automatically discovers mappers that surpass human expert designs in scientific applications by up to 1.34×1.34$\times$1.34 × speedup.
-For parallel matrix multiplication algorithms, our mapper achieves up to 1.31×1.31$\times$1.31 × of the expert-designed solution.
-To achieve this, we simplify the complexity of low-level code generation by introducing a domain-specific language (DSL) that abstracts the low-level system programming details and defines a structured search space for LLMs to explore.
-To maximize the application performance, we use an LLM optimizer to improve an agentic system that generates the mapper code.
-As a result, this approach significantly reduces the workload for performance engineers while achieving substantial performance gains across diverse applications.
-Finally, our results demonstrate the effectiveness of LLM-based optimization in system design and suggest its potential for addressing other complex system challenges.
+Modern scientific discovery increasingly relies on high-performance computing for complex modeling and simulation.
+A key challenge in improving parallel program performance is efficiently mapping tasks to processors and data to memory, a process dictated by intricate, low-level system code known as mappers.
+Developing high-performance mappers demands days of manual tuning, posing a significant barrier for domain scientists without systems expertise.
+We introduce a framework that automates mapper development with generative optimization, leveraging richer feedback beyond scalar performance metrics.
+Our approach features the Agent-System Interface, which includes a Domain-Specific Language (DSL) to abstract away low-level complexity of system code and define a structured search space, as well as AutoGuide, a mechanism that interprets raw execution output into actionable feedback.
+Unlike traditional reinforcement learning methods such as OpenTuner, which rely solely on scalar feedback, our method finds superior mappers in far fewer iterations.
+With just 10 iterations, it outperforms OpenTuner even after 1000 iterations, achieving 3.8×3.8$\times$3.8 × faster performance.
+Our approach finds mappers that surpass expert-written mappers by up to 1.34×1.34$\times$1.34 × speedup across nine benchmarks while reducing tuning time from days to minutes.
+Machine Learning, ICML
 1 Introduction
-Task-based programming (Slaughter et al., 2015; Bauer et al., 2012; Augonnet et al., 2009; Chamberlain et al., 2007; Kaiser et al., 2014; Heller et al., 2017; Chandra et al., 2001; Duran et al., 2011; Moritz et al., 2018; Barham et al., 2022) has recently emerged as a prominent paradigm in parallel programming.
-The core idea is to decompose computations into self-contained functions called tasks that do not communicate with other tasks except through their arguments (typically tensors or multi-dimensional arrays).
-A performance-critical aspect of executing task-based applications is mapping, which involves assigning tasks to processors, mapping data to physical memories, and managing other low-level hardware resources.
-A concrete mapping policy, implemented as code, is referred to as a mapper.
-Mappers focus solely on improving performance without altering the correctness of the application’s output.
-The performance impact of a good versus a poor mapper can be substantial, often resulting in differences of an order of magnitude (Abdel-Gawad et al., 2014; Galvez et al., 2017).
-Currently, writing mappers is a manual, labor-intensive, and time-consuming process that requires deep knowledge of the application, the machine, and the low-level system’s C++ mapping APIs.
-Even for experienced engineers, developing a high-performance mapper can take several days, requiring multiple rounds of refinement (SFX Teixeira et al., 2023).
-In this paper, we focus on using LLMs to automatically generate mappers.
-There are two main challenges in using LLMs for mapper code generation.
-First, given a specific mapping policy described in natural language, LLMs struggle to generate the corresponding mapper code, which consists of a few hundred lines of low-level C++ code.
-This code generation task is significantly different from function-level code generation, which is more isolated and modular, where LLMs have demonstrated strong performance (Chen et al., 2021; Austin et al., 2021).
-In our case, generating C++ mappers requires a deep understanding of low-level systems and broader contextual dependencies in large-scale system code, which is beyond the capabilities of current LLMs (Jimenez et al., 2023; Du et al., 2023; Wang et al., 2024).
-The second challenge lies in identifying the optimal mapping strategy to maximize application performance.
-Many design problems in systems are naturally constrained optimization problems.
-Generating mappers that are syntactically correct is insufficient.
-To address the first challenge, we design a Domain-Specific Language (DSL) for mapper generation.
-The DSL provides a high-level interface that encapsulates all the critical decisions a mapper has to make without interacting with the intricate low-level C++ APIs.
-The DSL defines a search space in its abstractions, which breaks a monolithic code generation down to generating multiple independent, modular components, which is much easier for code generation due to less contextual dependence between components.
-For the second challenge, oirst, structuring the search space through the DSL’s decomposition and modularization of the mapper, and second, applying reinforcement learnformulating the search problem as an optimization problem where the objective is to maximize the performance of the parallel program.
-Similar approaches have achieved significant results in discovering faster algorithms for matrix multiplication (AlphaTensor) (Fawzi et al., 2022) and generating mathematical proofs (AlphaGeometry) (Trinh et al., 2024), both require searching through a discrete space for an optimal solution.
-Recently, a great deal of work has demonstrated that instruction-tuned LLMs have an inherent ability to “optimize” for an objective, such as minimizing a function (Nie et al., 2024; Cheng et al., 2024), prompt optimization (Yang et al., ; Khattab et al., 2023), and molecule design optimization (Yuksekgonul et al., 2024).
-We propose a fully automated agent-based system to generate mappers and optimize for parallel application performance, illustrated in Figure 1.
-Our solution allows LLMs to refine their mapping strategies efficiently, reducing the time required to develop high-performance mappers from days to just a few minutes.
-Figure 1: We illustrate an agent-based optimization for high-performance mapper code generation, where the mapper is written in a Domain-Specific Language (DSL) and iteratively refined by an agent through performance feedback.
-Initially, all tasks are mapped to the CPU and system memory.
-The LLM then improves by mapping computation-intensive tasks to the GPU and GPU’s FrameBuffer memory to minimize task execution time.
-Finally, tasks that require significant data exchange between the CPU and GPU are mapped to the ZeroCopy memory, enabling shared access and minimizing data movement overhead.
-Our experiments show that mappers generated by LLMs not only match but can surpass expert-written mappers, achieving up to 1.34×1.34$\times$1.34 × speedup across 9 benchmarks.
-Notably, the search process enhances the performance of parallel matrix multiplication algorithms by up to 1.31×1.31$\times$1.31 ×.
-We also demonstrate, through ablation studies, the critical role of high-quality feedback in guiding LLM optimizers to discover more efficient mapping strategies.
+Modern scientific discovery depends on advanced software tools for modeling and simulation (Stocks et al., 2024; Wang et al., 2024; Ltaief et al., 2024).
+Computational scientists, including physicists, chemists, and biologists, rely on high-performance computing to tackle complex problems.
+These scientific computations dominate workloads on the world’s most powerful supercomputers (Exascale Computing, ).
+However, many domain scientists lack expertise in computer science, and therefore having difficulties in optimizing their programs because of the complexity and scale of the underlying machines.
+Even for experts, finding and fixing performance problems resulting from program modifications or when porting to a new machine is often time-consuming.
+Any progress on automating performance tuning is of great benefit in this domain.
+Task-based programming (Slaughter et al., 2015; Bauer et al., 2012; Augonnet et al., 2009; Chamberlain et al., 2007; Moritz et al., 2018; Barham et al., 2022) has emerged as a promising approach to high performance computing.
+The paradigm involves decomposing computations into independent tasks that communicate exclusively through their arguments.
+A key advantage of task-based systems is that the performance tuning problem is factored out into a separate mapping: an assignment of tasks to processors and data to particular memories.
+High-quality mapping, achieved through a well-designed mapper (implemented as code), can significantly improve performance, often by an order of magnitude (Galvez et al., 2017).
+However, currently writing mappers remains a labor-intensive process, as it requires deep knowledge of applications, hardware, and low-level system APIs.
+In addition, this process is highly application-specific, input-specific, and machine-specific, often taking experts several days of meticulous tuning to achieve high performance.
+This challenge is especially pronounced for domain scientists, who typically lack the necessary expertise in computer systems and code optimization.
+Automating mapper development would enable scientists to focus on their own domain of expertise while fully utilizing the capabilities of high-performance computing systems.
+Figure 1: Iterative mapper refinement with agent-based generative optimization.
+The system leverages the Agent-System Interface, which consists of the Domain-Specific Language (DSL) and AutoGuide.
+The DSL abstracts away the low-level system code, defining a search space for mapping strategies, while AutoGuide interprets execution results into actionable guidance.
+As iterations progress, the mapper evolves to improve performance.
+In this paper, we introduce a system powered by large language models (LLMs) to automate both the generation and optimization of mapper code.
+The first challenge stems from the complexity of generating mapper code due to the original low-level programming system, which exposes the agent to intricate system APIs, coupled with the problem that raw feedback messages from the system are often uninformative to the agent.
+The second challenge involves optimizing mapper performance.
+Specifically, it consists of (1) defining an appropriate search space and (2) devising efficient methods to find optimal mappers, thereby maximizing parallel program performance.
+To address the first challenge, we propose an Agent-System Interface (ASI), as shown in Figure 1, an abstraction layer between the agent and the system that simplifies code generation and provides more meaningful feedback to the agent.
+At the core of ASI is a Domain-Specific Language (DSL), a high-level interface that encapsulates all performance-critical decisions required to generate a mapper.
+The DSL abstracts away the complexity of low-level system code with a compiler.
+Additionally, the DSL defines a structured search space, enabling systematic exploration of mapping strategies.
+We also design and implement the AutoGuide mechanism to interpret raw execution output into informative and actionable guidance.
+This mechanism allows the agent to iteratively optimize the mapper by leveraging enriched feedback to update its strategy.
+For the second challenge, we adopt the generative optimization approach, a recent advance in optimization techniques.
+Unlike traditional methods such as reinforcement learning (Ansel et al., 2014), which rely solely on scalar rewards, generative optimization can utilize richer forms of feedback, such as error explanations and actionable suggestions expressed in natural language.
+This agentic optimization workflow has previously proven to be effective across various domains (Nie et al., 2024; Cheng et al., 2024; Yang et al., ; Khattab et al., 2023; Yuksekgonul et al., 2024).
+Our work is the first to apply such technique to the domain of system optimization.
+Our experiments demonstrate that mappers optimized by LLM-powered agents not only match but often surpass expert-written mappers, achieving up to 1.34×1.34$\times$1.34 × speedup across nine benchmarks.
+Since expert-written mappers set the highest standard, surpassing them is a notable accomplishment.
+At the same time, our method significantly reduces mapper tuning time from days to minutes, making high-performance mapping more accessible to domain scientists.
+To further highlight the advantage of generative optimization, we compare it against OpenTuner, a reinforcement learning-based autotuning framework.
+Our generative optimizer finds mappers $\times1$ faster than OpenTuner when both run for 10 iterations and still maintains a 3.8×3.8$\times$3.8 × advantage even when OpenTuner runs for 1000 iterations.
+Furthermore, ablation studies underscore the necessity of the agent-system interface design in achieving these performance gains.
 Our contributions are as follows:
   1. 1.
-    Development of a Domain-Specific Language (DSL) to Simplify Mapper Generation: We design a DSL to address the complexities of direct code generation in low-level systems.
-This DSL provides a higher-level abstraction, enabling LLMs to generate mapping code more effectively by encapsulating the complexities of low-level C++ APIs.
+    Design of an Agent-System Interface: We introduce an abstraction layer that simplifies mapper code generation and provides guidance to the agent.
+The Domain-Specific Language (DSL) defines a search space, allowing the agent to explore mapping strategies without dealing with low-level system code.
+AutoGuide interprets raw execution output into targeted feedback, enabling the agent to refine mapper code more effectively.
   2. 2.
-    Formulation of Mapper Generation as a Discrete Optimization Problem: We formulate the task of mapper code generation as a discrete optimization problem with the objective of maximizing application performance.
-Our DSL defines a more structured and constrained search space for LLM to explore.
+    Generative Optimization for Systems: We introduce generative optimization to improve system performance, leveraging richer feedback such as error messages and actionable suggestions in natural language.
+Unlike reinforcement learning methods like OpenTuner, which rely solely on scalar feedback, our method identifies better mappers in far fewer iterations.
+With only 10 iterations, it outperforms OpenTuner by 3.8×3.8$\times$3.8 × even after 1000 iterations.
   3. 3.
-    Experimental Validation and Performance Comparison: Our DSL significantly improves the success rate of mapper code generation for a given mapping strategy from 0% in C++ to 80% in our DSL.
-Leveraging high-quality feedback, our agent-based solution discovers mappers that achieve up to 1.34×1.34$\times$1.34 × speedup compared with expert-designed mappers, enhancing the performance of parallel matrix multiplication algorithms by up to 1.31×1.31$\times$1.31 ×.
+    Empirical Evaluation of Performance: Our agent-based solution achieves up to 1.34×1.34$\times$1.34 × speedup across nine benchmarks, surpassing expert-written mappers while reducing tuning time from days to minutes.
+We highlight the critical role of the agent-system interface through ablation studies, demonstrating its impact on achieving the performance gains.
 2 Related Work
-Mapping in Parallel Programming Many parallel programming systems allow users to make their own mapping decisions, such as Legion (Bauer et al., 2012), StarPU (Augonnet et al., 2009; 2010), Chapel (Chamberlain et al., 2007), HPX (Kaiser et al., 2014; Heller et al., 2017), Sequoia (Fatahalian et al., 2006), Ray (Moritz et al., 2018), and Pathways (Barham et al., 2022).
-Several techniques have been proposed to automate mapping, including machine learning models (O’Boyle et al., 2013; Wang & O’Boyle, 2009), static analysis (Poesia et al., 2017; Ren et al., 2008), and auto-tuning (SFX Teixeira et al., 2023).
-Unlike previous work, we use an agent-based approach with LLMs, exploring a larger search space for mappers than traditional methods.
-Automated Code Generation With the rise of LLMs, various models (Chen et al., 2021; Nijkamp et al., 2022; Li et al., 2023; Ouyang et al., 2022; Wei et al., 2023) have been trained on code, enabling them to assist with programming tasks and automate software development (Shypula et al., 2023; Yang et al., 2024b; Jain et al., 2024).
-While LLMs have shown success in generating isolated function-level Python code (Chen et al., 2021; Austin et al., 2021), they face challenges when generating system-level C++ code in large repositories due to complex APIs and broader contextual dependencies (Du et al., 2023; Wang et al., 2024).
-Thus, adopting LLMs to generate code in real-world software remains difficult (Jimenez et al., 2023).
-Our work addresses this by developing a domain-specific language (DSL) that provides a high-level abstraction to help LLMs generate mapper code more effectively.
-Automated System Design Optimization The use of machine learning for optimizing system design has garnered significant attention in recent years.
-Learned cost models have proven effective in optimizing deep learning operators on GPUs (Zheng et al., 2020a; b; 2022b).
-Methods in reinforcement learning and multi-armed bandit have been applied to tasks such as chip floorplanning (Mirhoseini et al., 2021) and program autotuning (Ansel et al., 2014), automatic vectorization of programs (Haj-Ali et al., 2020a), and optimizing compiler phase ordering (Haj-Ali et al., 2020b).
-Previous work primarily focuses on hyperparameter configurations, while our work aims to improve system performance by directly generating mapper code.
-Optimization with LLMs
-Recently, LLMs have been increasingly applied to optimization problems traditionally found in numerical domains (Boyd & Vandenberghe, 2004).
-AhmadiTeshnizi et al.
-(2023) addressed structured optimization challenges, such as mixed-integer linear programming (LP), while Yang et al.
-(2024a) applied LLMs to unstructured problems, treating them as black-box optimizations.
-(Nie et al., 2024) highlighted the role of feedback in efficiently finding global minima for black-box functions.
-Cheng et al.
-(2024) formally defined optimization problems solvable by LLMs and proposed an optimizer based on execution graphs.
-More recently, Yuksekgonul et al.
-(2024) introduced an LLM optimizer inspired by gradient descent.
-These advancements demonstrate the potential of LLMs to optimize beyond task completion, improving performance metrics across various domains.
-3 Mapper Generation Task
-Task Definition
-The problem we address is the automated generation of high-performance mappers for the Legion parallel programming framework (Bauer et al., 2012).
-Mappers determine how tasks are assigned to processors and how data is placed in memory to optimize performance.
-A well-designed mapper can achieve up to 10$\times$ speedup compared to random mapping strategies.
-However, mappers are highly context-specific and must be carefully tailored to an application’s input and the machine’s architecture.
-Finding an optimal mapper is akin to solving a combinatorial optimization problem—a process that is time-consuming, labor-intensive, and traditionally performed by experienced performance engineers.
-The search space for discovering the best mapping strategy is vast, growing to $2^{14}$ even for the simplest scientific applications (as shown in prior work with a smaller search space than ours (SFX Teixeira et al., 2023)), making it challenging to efficiently explore all possible solutions.
-Moreover, even with a clear mapping strategy, writing the corresponding mapper requires experts to produce hundreds of lines of low-level C++ code, a task that can take several hours.
-As a result, the entire performance tuning process can take several days to complete.
-Figure 2: Mappers decide the placement of each task in the task graph to processors, the placement of data to memory, and how the iteration space of data is partitioned and mapped to different processors.
-Mapping Decisions
-Now we elaborate on the decisions that a mapper has to make.
-The first critical decision is the processor selection for each task, determining whether a task is better suited for GPUs, CPUs, or OpenMP runtime.
-This choice depends on factors such as task size, GPU memory, and kernel launch latency.
-For example, tiny tasks that require very little computation may prefer to run on CPUs due to the GPU kernel launch overhead, whereas tasks with large memory footprints might prefer to be assigned to OpenMP or CPU when GPU memory is insufficient.
-Next, the memory placement of data across different memory spaces is crucial to performance.
-A mapper must decide where to place data — in the GPU’s FrameBuffer for faster access, in ZeroCopy memory for shared access between CPU and GPU, or in CPU system memory for large data.
-Each choice introduces a trade-off between memory access speed, memory usage, and transfer overhead.
-The memory layout option determines the optimal memory arrangement for data structures.
-Depending on the memory access patterns and the underlying hardware, selecting between a Struct of Arrays (SOA) or an Array of Structures (AOS) layout, along with constraints on array ordering (Fortran-order or C-order) and memory alignment, can significantly impact performance due to cache efficiency and data locality.
-Such choice is task-dependent and processor-dependent.
-Finally, mappers need to decide how to perform index mapping.
-As shown in Figure 2, index mapping controls how the data (or more accurately, parallel for loops of task launches) is partitioned and mapped to the processors.
-The data (or tensors) can be multi-dimensional, and the distributed machines can also be viewed as a processor space (nodes and processors within the node).
-Index mapping controls the mapping between the two index spaces.
-Prior work (Unger et al., 2022; Zheng et al., 2022a) shows that searching over how data is partitioned and mapped can significantly change inter-processor communication volume, which affects performance.
-Commonly-used index mapping functions are shown in Section A.3.
-4 A New Approach: Generation and Optimization in a DSL
-4.1 Design a Domain-Specific Language
-We designed a domain-specific language for mapper generation.
-We distill the key performance-critical aspects of mapping into language constructs that are easily expressed in the DSL.
-By providing a higher-level abstraction than C++ APIs, the DSL simplifies interfacing with LLMs.
-Also, such a modular design enables LLMs to efficiently address the complex optimization challenges of generating high-performance mappers.
-To implement our DSL, we develop a compiler that can translate the mapper written in DSL into low-level C++ mapping APIs.
-Figure 3(a) presents an example DSL mapper that illustrates the key features of our domain-specific mapping language.
-By contrast, Figure 3(b) shows the (already simplified) code snippet extracted from a C++ mapper.
-Notably, the C++ snippet covers only part of the functionality that the IndexMap statement (which we discuss in detail below) in the DSL provides.
-While a typical C++ mapper requires approximately 400 lines of code, the corresponding DSL mapper can achieve the same functionality in just about 30 lines on average — reflecting a 14×14$\times$14 × reduction, as detailed in Table 1 (based on the 9 applications used in Section 5.2 and Section 5.3).
-This substantial reduction in code complexity makes the DSL a more accessible target for LLM-based code generation, as it abstracts away many low-level implementation details.
+Mapping in Parallel Programming Many parallel programming systems allow users to make their own mapping decisions, such as Legion (Bauer et al., 2012), StarPU (Augonnet et al., 2009, 2010), Chapel (Chamberlain et al., 2007), HPX (Kaiser et al., 2014; Heller et al., 2017), Sequoia (Fatahalian et al., 2006), Ray (Moritz et al., 2018), TaskFlow (Huang et al., 2021), and Pathways (Barham et al., 2022).
+Several techniques have been proposed to automate mapping, including machine learning models (O’Boyle et al., 2013; Wang & O’Boyle, 2009), static analysis (Poesia et al., 2017; Ren et al., 2008), reinforcement learning (Ansel et al., 2014; Mirhoseini et al., 2017) and auto-tuning (SFX Teixeira et al., 2023).
+We use an agent-based approach with LLMs and explore a larger search space for mappers than traditional methods.
+Agentic Frameworks Agents powered by Large Language Models (LLMs) play a critical role in decision-making, planning, tool integration, and solving complex problems in dynamic environments (Guo et al., 2024).
+Many agentic frameworks have been developed (Yao et al., 2022; Wu et al., 2023; Li et al., 2023; Hong et al., 2023), with uses spanning domains such as software engineering (Gur et al., 2023; Yang et al., 2024b; Jin et al., 2024), robotics (Kannan et al., 2024), healthcare (Li et al., 2024), education (Ramirez & Esparrell, 2024), and knowledge engineering (Shao et al., 2024).
+Our work is the first to apply an agentic workflow to iteratively optimize mapper code, improving the performance of parallel programs.
+AI for Systems The application of AI to optimize system design has gained significant traction in recent years.
+Techniques such as deep learning (Zheng et al., 2020a, b, 2022b) and gradient-boosted trees (Feng et al., 2023) have been used to predict program execution times for performance optimization.
+Reinforcement learning methods have addressed challenges in chip floorplanning (Mirhoseini et al., 2021), autotuning (Ansel et al., 2014), auto-vectorization (Haj-Ali et al., 2020a), and compiler phase ordering (Haj-Ali et al., 2020b).
+While previous efforts have predominantly relied on traditional approaches for cost prediction and optimization, our work uses the recent advances in generative optimization to tackle complex system challenges.
+Generative Optimization
+Recent work has explored the use of LLMs for optimization problems traditionally tackled with numerical methods, including mixed-integer programming (AhmadiTeshnizi et al., 2024a, b) and numerical optimization (Nie et al., 2024).
+A key advantage of generative optimization is its ability to iteratively refine solutions using diverse forms of feedback.
+For example, Cheng et al.
+(2024) applies generative optimization to robotic manipulation and game playing, while Yuksekgonul et al.
+(2024) optimizes prompts and molecular designs.
+While reinforcement learning has been applied to system optimization, the potential of LLM-driven optimization in systems remains unexplored.
+Our work explores whether generative optimization with richer feedback outperforms traditional methods using scalar rewards in system optimization.
+3 Problem Definition
+Motivation and Challenges
+The concrete problem we address is the automated generation of high-performance mappers for the Legion parallel programming framework (Bauer et al., 2012).
+Mappers dictate task scheduling and data placement.
+A well-designed mapper can achieve orders-of-magnitude speedup over naive strategies.
+However, automating mapper generation is challenging due to two key factors.
+First, the complexity of low-level system code.
+Implementing a mapper requires writing hundreds of lines of intricate C++ code, demanding expertise in system internals.
+Second, the vast search space of mapping strategies.
+The search space grows exponentially with the number of tasks and arguments.
+Search Space and Performance Impact
+As illustrated in Figure A1, the search space of mappers involves multiple decisions, each influencing performance.
+The first key aspect is processor selection, which determines whether a task runs on GPUs, CPUs, or the OpenMP runtime.
+This choice depends on factors such as task size, GPU memory capacity, and kernel launch overhead.
+For instance, small tasks may prefer CPUs due to the overhead of launching GPU kernels, while tasks with large memory footprints may run on CPUs when GPU memory is insufficient.
+Another crucial dimension is memory placement, which dictates where data is stored.
+A mapper must decide whether to place data in the GPU’s FrameBuffer for fast access, ZeroCopy memory for CPU-GPU sharing, or CPU system memory for more available storage.
+Each option presents trade-offs between access speed, memory usage, and data transfer overhead.
+Additionally, memory layout further expands the search space, with decisions on Struct of Arrays (SOA) vs.
+Array of Structures (AOS), data ordering (Fortran-order vs.
+C-order), and alignment constraints (e.g., 128-byte alignment) significantly affecting cache efficiency and performance.
+Finally, an important idiom in high-performance computing is launching tasks over partitioned data.
+Index mapping determines how data partitions and task executions are distributed across multiple processors.
+For consistency, we can represent data partitioning as a tensor of data partitions, the machine as a tensor of processors, and tasks operating on the partitioned data as a tensor of tasks.
+The way data and task indices are mapped to processor indices affects inter-processor communication, a key factor in performance (Unger et al., 2022; Zheng et al., 2022a).
+4 Our Approach: Agent-System Interface
 1# Map task0 to GPU.
@@
 19IndexTaskMap task4 cyclic
-(a) An example mapper in DSL
+(a) An example mapper in Domain-Specific Language (DSL)
 1void slice_task(const Task& task,
@@
 9 Rect<2> blocks(zeroes, num_blocks - ones);
-10 ...// **126 lines of C++ code ommitted**
+10 ... // 126 lines of C++ code omitted here
 11 for (PointInRectIterator<2> it(blocks); it() != NULL; it++)
+12 {
 13 DomainT<2,coord_t> slice_space;
@@
 18 }
-18}
-(b) Simplified code snippet from a C++ mapper
-Figure 3: A DSL mapper and a short code snippet from C++ mapper.
-The C++ mapper has already been simplified, and the shown C++ snippet only contributes to part of the functionality achieved by the IndexTaskMap statement on Line 19 in the DSL mapper.
-Next, we explain the DSL’s design, highlighting the immense search space it opens for optimization, which makes generating high-performance mappers a challenging task.
-The grammar of the DSL is shown in Section A.1.
-The Task statement (Line 2) performs processor selection for tasks.
-Although it may seem straightforward to always prefer GPUs, this decision is complicated by other factors such as limited GPU memory and kernel execution time as explained in Section 3.
-This is a per-task decision.
-The Region statement (Line 5) performs memory placement for data.
-Possible choices include GPU’s FrameBuffer or ZeroCopy memory, CPU’s System memory or RDMA memory.
-Such decisions need to be made for each argument of each task, which opens up a huge trade-off space between memory usage, task execution time, and data transfer costs as explained in Section 3.
-The Layout statement (Line 9) defines memory layouts, supporting both SOA (Struct of Arrays) and AOS (Array of Structures), as well as array ordering constraints like C_order, F_order, and memory alignment.
-Optimizing memory layout is crucial for performance, as access patterns vary with tasks, processors, and data structures.
+19}
+(b) Code snippet from a C++ mapper
+Figure 2: Comparison of a DSL mapper and a C++ mapper.
+The DSL’s declarative, high-level design abstracts away the complexity of low-level C++ code, serving as the core of the Agent-System Interface.
+The highlighted boxes illustrate how the same functionality, which requires extensive C++ system code, can be expressed concisely in just a few lines in DSL.
+4.1 Domain-Specific Language Design
+A key challenge in automating mapper generation with a coding agent is the complexity of low-level system code, which requires intricate C++ implementations.
+To address this, we design a high-level Domain-Specific Language (DSL) as the core of our Agent-System Interface (ASI).
+The DSL provides a structured search space for mapping strategies while abstracting away low-level implementation details.
+Unlike C++, which demands imperative specifications of mapping policies, our DSL adopts a declarative design, allowing users to specify what to achieve rather than how to implement it.
+Most critically, the DSL separates concerns, enabling multiple aspects of mapping decisions to be expressed independently rather than being entangled in low-level system APIs.
+This design reduces code complexity and naturally provides a search space for the agent to explore.
+To implement it, we develop a compiler that translates DSL into the low-level C++ APIs.
+As illustrated in Figure 2, the complexity of DSL code is significantly lower than that of C++.
+Figure 2(a) provides an example of a DSL mapper, highlighting the key features of our DSL.
+In contrast, Figure 2(b) shows a snippet from a C++ mapper, emphasizing the intricacy of low-level implementation details.
+According to Table A1, using the DSL results in an average lines of code reduction of 14×14$\times$14 ×.
+This substantial reduction makes DSL a more suitable target for LLM code generation, as it abstracts away the complexities inherent in low-level systems.
+As we will show in Section 5.2, LLMs generate DSL code more effectively, despite DSL having no examples in LLM training corpora, whereas C++ is widely represented.
+Next, we describe the DSL’s design, emphasizing its declarative nature and structured search space. Section 3 details the performance impact of each decision.
+The Task statement (Line 2) defines processor selection for each task, choosing between CPU, GPU, or OpenMP.
+Line 2 specifies that instances of task0 should run on GPUs.
+This decision is made per task; note that the search space expands exponentially with the number of tasks.
+The Region statement (Line 5) controls memory placement for data arguments.
+Line 5 specifies that all tasks using ghost_region should place the data in GPU ZeroCopy memory.
+Other choices include GPU FrameBuffer memory and CPU System Memory.
+This decision is made per task and per argument, causing the search space to grow exponentially.
+The Layout statement (Line 9) defines memory layouts.
+Line 9 enforces a C_order axis ordering, an SOA layout, and a 64-byte memory alignment for all data used by all tasks mapped to all processors.
+Alternative choices include F_order, AOS, and various alignment strategies.
 This is a per-task, per-data, per-processor decision.
-The IndexTaskMap statement (Line 19) enables the mapping of tasks to processors using custom functions, such as cyclic or block mapping strategies.
-This creates a mapping between the index space defined in application code (e.g., for loops) and the processor space of distributed machines.
-Many parallel operators (e.g., matrix multiplications) are based on such loops, and deciding how to map them to distributed machines is inherently difficult.
-The DSL enables users to express arbitrary arithmetic mappings between two index spaces, significantly increasing the complexity of the search space.
-Additionally, we introduce primitives that transform the processor space, allowing for more flexible and precise index mapping decisions, as detailed in Section A.2.
+The IndexTaskMap statement (Line 19) controls index mapping using a customized function.
+Line 12 defines the mapping function that establishes the correspondence between two index spaces: the task index space (represented by task.ipoint) defined in the application code (e.g., for loops) and the processor space of the distributed machine (represented by Machine(GPU)).
+The DSL allows users to express arbitrary arithmetic mappings between the two index spaces.
+This decision applies to each task group launched by parallel for loops.
+4.2 Generative Optimization via AutoGuide
+Figure 3: Agent Optimization Process.
+The mapper agent takes server specifications and application-specific information as input, generates mapper code, and executes it alongside the application on the server.
+Raw execution feedback is enriched using the AutoGuide mechanism and iteratively refined by an LLM optimizer to improve performance.
+Case Raw Execution Output AutoGuide
+Explain Suggest
+Case 1 Execution Error: Assertion failed: stride does not match expected value. Memory layout is unexpected. Adjust the layout constraints or move tasks to different processor types.
+Case 2 Performance Metric: Execution time is 0.03s. N/A Move more tasks to GPU to reduce execution time.
+Table 1: AutoGuide Feedback Mechanism.
+The AutoGuide mechanism interprets raw execution output from the runtime system, providing more informative error explanations and suggestions for mapper modifications.
+It is implemented via keyword matching.
+Additional examples are shown in Table A2.
+We formulate mapper generation as an online optimization problem.
+Given a triplet $(\Theta,\omega,\mathcal{T})$, where $\Theta$ is a set of possible mappers, $\omega$ is an optimization objective, and $\mathcal{T}$ is a function that takes a mapper $\theta\in\Theta$ as input, $(f,g)=\mathcal{T}(\theta)$ and returns $f$, the feedback from executing the mapper (i.e., the measured performance after running the application code with the generated mapper), and $g$, the process graph tracing how the mapper was generated.
+In our setup, mapper performance is deterministic, as we carefully control all sources of randomness in the environment.
+If the parameter space were numerical, this online optimization problem could be addressed using bandit algorithms (Lattimore & Szepesvári, 2020), reinforcement learning (Sutton & Barto, 2018), or Bayesian optimization (Snoek et al., 2012), but these methods are less efficient when the parameter search space is large and discrete (i.e., text).
+In this online optimization problem, we leverage the DSL to structure the parameter space to improve the efficiency of optimization.
+Here, $\theta$ represents the program code, while $\omega$ and $f$ are expressed as text.
+We adopt generative optimization, leveraging LLMs as optimizers given the objective in text form.
+This emergent optimization behavior has been recently observed and applied across various domains (Yang et al., 2024a; Cheng et al., 2024; Yuksekgonul et al., 2024; Patel et al., 2024).
+Optimization Process
+We present the optimization process in Figure 3.
+The agent takes two inputs: server specifications (e.g., CPU/GPU counts) and application information (e.g., task lists, data arguments).
+It generates mapper code, which is executed alongside the application code on the server.
+Raw execution feedback from the runtime is augmented with the AutoGuide mechanism and fed back to the LLM, iteratively refining the agent for improved mapper code generation.
+Coding Agent
+Our mapper agent improves mapping decisions by iteratively generating DSL code.
+A high-level schema of the mapper agent is shown in Figure 3.
+The mapper agent is implemented as a Python program in the Trace (Cheng et al., 2024) framework, where we decompose the task of generating a monolithic mapper into independent code segments.
+This decomposition allows the agent to decide what code to generate for each segment separately.
+This approach is effective because our DSL design eliminates unnecessary dependencies between mapping decisions.
+Our modularization strategy aligns with least-to-most prompting (Zhou et al., 2022).
+AutoGuide
+The AutoGuide feedback mechanism is designed based on three key motivations: (1) generative optimization benefits from natural language feedback rather than relying solely on scalar values, (2) raw execution output from the runtime system is often too uninformative to effectively guide the agent’s decisions, and (3) domain heuristics known to systems researchers can be naturally expressed in language (e.g., most tasks run faster on GPUs than CPUs).
+To address these needs, AutoGuide helps the agent by explaining opaque error messages and suggesting mapper modifications.
+As shown in Table 1, it interprets uninformative execution output into actionable insights, with additional examples in Section A.4.
+The implementation relies on keyword matching over the raw execution output.
+An ablation study in Section 5.3 demonstrates its effectiveness in our experiments.
+5 Evaluation
+Experiments are conducted on one node with two Intel 10-core E5-2640 v4 CPUs, 256G main memory, and four NVIDIA Tesla P100 GPUs. We use gpt-4o-2024-08-06.
+5.1 Speedup of Application Performance
+Figure 4: Performance Comparison.
+Normalized throughput for 9 benchmarks, comparing expert mappers, random mappers, the average optimization trajectories of Trace, OPRO, and OpenTuner in 10 iterations across 5 runs, and the best mappers found by Trace.
+Figure 5: Comparison of Trace (generative optimizer) and OpenTuner (traditional RL) over 1K iterations (averaged across all 9 benchmarks)
+We evaluate our approach using 9 benchmarks.
+Circuit (Bauer et al., 2012) is a simulation benchmark that models electrical circuit behavior by simulating currents and voltages across interconnected nodes and wires.
+Stencil (Van der Wijngaart & Mattson, 2014) simulates a 2D grid where each point’s value is updated based on a stencil pattern determined by its neighbors.
+Pennant (Ferenbaugh, 2015) models unstructured mesh Lagrangian staggered-grid hydrodynamics, commonly used for simulating compressible flow.
+The remaining six benchmarks – Cannon’s (Cannon, 1969), SUMMA (Van De Geijn & Watts, 1997), PUMMA (Choi et al., 1994), Johnson’s (Agarwal et al., 1995), Solomonik’s (Solomonik & Demmel, 2011), and COSMA (Kwasniewski et al., 2019) – are widely studied parallel matrix multiplication algorithms, which we discuss in more detail in Section A.3.
+In this experiment, we evaluate the performance of the mappers with the following baselines.
+Expert-Written Mappers.
+These mappers are manually developed by domain scientists who spend years mastering computational science.
+Writing mappers in parallel programming frameworks is another challenge, and tuning them for specific applications can take days.
+Randomly Generated Mappers. These mappers were randomly generated with 10 different random seeds, sampling from the entire search space of each application. We report the average performance.
+Agent-Optimized Mappers.
+Using Trace (Cheng et al., 2024), we evaluated the Trace and OPRO (Yang et al., ) search algorithms, running 10 iterations per application.
+To account for stochastic output, we repeated the process 5 times and report the average.
+The best mapper from Trace across runs is also reported.
+OpenTuner Mappers.
+OpenTuner (Ansel et al., 2014) is a program autotuning framework that uses reinforcement learning to optimize performance based on scalar feedback.
+We provided execution time as feedback, with a high penalty for failures.
+Results
+We use normalized throughput as the performance metric in Figure 4, where higher values indicate better performance.
+The throughput is normalized relative to the expert-written mappers.
+All the best mappers found by Trace can match or surpass the expert-written mappers, underscoring the effectiveness of agent-based generative optimizer.
+Random mappers consistently exhibit low performance across all applications, emphasizing the critical role of mapping decisions.
+When comparing optimization trajectories, Trace performs similarly to OPRO, and significantly outperforms OpenTuner.
+To further compare the agent-based optimizer with traditional reinforcement learning, we extended OpenTuner’s optimization iterations from 10 to 1000, as shown in Figure 5, where the x-axis is the log-scale of iterations and the y-axis represents relative throughput (averaged across all 9 benchmarks).
+Notably, Trace achieves a 3.8×3.8$\times$3.8 × speedup over OpenTuner even when OpenTuner is run for 1000 iterations.
+When both are limited to 10 iterations, Trace outperforms OpenTuner by $\times1$, demonstrating its ability to quickly identify high-performance mappings.
+This highlights the superiority of Trace (generative optimizer) over OpenTuner (traditional reinforcement learning).
+Moreover, Trace completes the entire optimization process in just 10 minutes per application, reducing mapper development time from days to minutes.
+Case Analysis
+The largest performance gain achieved by Trace over the expert mapper is observed in Circuit, with a speedup of 1.34×1.34$\times$1.34 ×.
+This improvement is primarily due to memory placement: the best mapper allocates two data collections to GPU FrameBuffer memory, while the expert mapper places them in GPU ZeroCopy memory.
+Despite a slight increase in inter-GPU communication costs, Trace reduces task execution time due to faster memory access, resulting in higher overall performance.
+For matrix-multiplication algorithms, the greatest speedup is seen in COSMA, with Trace achieving a 1.31×1.31$\times$1.31 × speedup over the expert mapper.
+This is attributed to Trace’s more efficient index mapping functions, which reduce inter-GPU communication by better distributing partitioned submatrices across GPUs.
+For additional context, examples of Trace mappers are presented in Section A.7.
+Code Generation Target Mapping Strategy Success Rate
+1 2 3 4 5 6 7 8 9 10
+C++ (single trial) 0%
+DSL (single trial) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 80%
+C++ (iterative refine) 0%
+DSL (iterative refine) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 100%
+Table 2: Code Generation Success Rates.
+Success rates for generating code across 10 mapping strategies described in natural language.
+The test evaluates whether the generated code compiles and passes execution tests.
+Generating DSL code significantly outperforms generating C++ for both settings.
+Symbols indicate results: – fails to compile, ✗ compiles but fails the test, and $\checkmark$ passes the test.
+Figure 6: Comparison of different feedback designs.
+0-Shot and 5-Shot are baselines.
+Execution provides only the raw execution output as feedback.
+Explain provides additional explanations of execution errors.
+Suggest offers mapper modification suggestions.
+All feedback is automatically generated.
+5.2 Ablation Study of DSL for Code Generation
+In Section 5.1, we demonstrate the overall effectiveness of our approach.
+Here, we conduct an ablation study on the DSL, the core of the Agent-System Interface.
+Since successful generation is the foundation of optimization, this subsection focuses on how well the DSL helps LLMs generate correct mappers compared to C++, rather than directly optimizing performance.
+Experiment Setup
+We designed 10 mapping strategies, described in natural language, to evaluate whether LLMs can generate correct code in both the DSL and the original low-level C++.
+The strategies are detailed in Section A.6.
+To ensure a fair comparison, identical prompt materials (documentation, examples, and starting code) were provided for both the DSL and C++.
+Success rates are measured based on whether the generated code passes predefined test cases, with results reported for single trials and iterative refinement, where the LLM is allowed up to 10 iterations to improve the code using compiler feedback.
+The evaluation is conducted with the DSPy (Khattab et al., 2023) framework.
+Results
+Table 2 shows that DSL achieves significantly higher generation success rates than C++ in both the single-trial and iterative refinement settings.
+This demonstrates the effectiveness of DSL’s design in abstracting system complexity and providing a high-level interface that enables LLMs to tackle complex system challenges in code generation.
+Incorporating iterative refinement with compiler feedback further improves success rates, resolving three compilation errors in C++ and two in the DSL.
+However, the gap between DSL mappers and C++ mappers remains substantial.
+Notably, these results are striking given that the DSL is a low-resource language with no pre-training or fine-tuning data, while C++ code is widely present in LLM training corpora.
+Analysis
+LLMs perform better with the DSL for two reasons.
+First, the semantic gap between natural language and code is smaller with the DSL than with C++.
+For example, writing a mapper to “align all data to 64 bytes in memory and use Fortran ordering” requires one line Layout * * * Align==64 F_order in the DSL because of its declarative design.
+In contrast, the C++ mapping API requires a sequence of operations to enforce alignment and ordering, which widens the semantic gap.
+Second, the DSL reduces the amount of code.
+As shown in Table A1, LLMs achieve an average reduction of 14×14$\times$14 × in lines of code, simplifying code generation.
+These results underscore the importance of a high-level agent-system interface.
+5.3 Ablation Study of the AutoGuide Feedback
+The AutoGuide mechanism provides enriched feedback to the agentic optimizer. We compare with alternative feedback designs.
+Experiment Setup
+We compare the following baselines.
+0-shot and 5-shot have no feedback, allowing the LLM to generate once with either 0 or 5 examples provided.
+Execution only provides raw execution feedback, Explain offers additional explanations for execution errors, and Suggest offers mapper modification suggestions.
+The Trace trajectory shown in Figure 4 uses the full AutoGuide mode with all Execution+Explain+Suggest.
+As an ablation study, we evaluate 3 benchmarks.
+Results and Analysis
+Figure 6 demonstrates that the full feedback mechanism consistently outperforms all reduced feedback variants.
+The 0-shot and 5-shot results perform the worst, underscoring the importance of feedback-based iterative refinement.
+This highlights the value of an agentic workflow, showing that performance improvements are not solely driven by prompting the LLM but are a direct result of the iterative refinement inherent in the workflow design.
+6 Conclusion
+In this paper, we introduced a system that leverages LLMs to automate mapper generation and optimization.
+The Agent-System Interface (ASI) simplifies code generation with a Domain-Specific Language (DSL), which abstracts away the low-level complexity of system code, and enriches execution feedback through AutoGuide, which interprets raw execution output into actionable guidance.
+We adopted generative optimization, allowing LLMs to refine mappers using rich textual feedback beyond scalar metrics.
+Unlike RL-based methods like OpenTuner, which rely on numerical rewards, our approach incorporates error explanations and targeted suggestions, accelerating search efficiency.
+Experiments show that agent-generated mappers outperform expert-written ones, achieving up to 1.34×1.34$\times$1.34 × speedup across nine benchmarks.
+Our method, running only 10 iterations, maintains a 3.8×3.8$\times$3.8 × advantage over OpenTuner even after 1000 iterations.
+By reducing mapper development time from days to minutes, our approach benefits computational scientists and demonstrates the effectiveness of generative optimization in system design.
+Acknowledgements
+We thank Yuhui Zhang, Chenyang Yang, Genghan Zhang, Qizheng Zhang, Mohammad R.
+Fadiheh, Ed Chen, Mert Yuksekgonul, Jiaxin Ge, Xinyan Hu, Tingyu Zhu, Simon Guo, Anne Ouyang, and Adhish Ramkumar for the feedback.
+We thank Ching-An Cheng and Ken Liu for their discussions and feedback.
+Impact Statement
+This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+Appendix A Appendix
+A.1 Illustration for Mapping
+We show an illustration for mapping in Figure A1.
+Figure A1: Mappers decide the placement of each task in the task graph to processors, the placement of data to memory, and how the iteration space of data is partitioned and mapped to different processors.
+A.2 Tables for Lines of Code Comparison
+We show the lines of code comparison between DSL and C++ mappers in Table A1.
 Application 1 2 3 4 5 6 7 8 9 Avg.
@@
 LoC Reduction $\times2$ $\times2$ 24×24$\times$24 × 12×12$\times$12 × 12×12$\times$12 × $\times1$ 13×13$\times$13 × $\times1$ 14×14$\times$14 × 14×14$\times$14 ×
-Table 1: Lines of Code (LoC) of mapper written in DSL, and LoC reduction compared with C++
-4.2 Learning to Generate via Interactive Feedback
-Figure 4: We show the high-level design of a mapper agent.
-The agent takes two inputs: server-related specifications and application-related information.
-When a mapper is generated by the agent, we run the mapper code on the server and get performance feedback.
-Then, we use an LLM to update each code block to improve the performance.
-1class MapperAgent(trace.Module):
-2
-3 @trace.bundle(trainable=True)
-4 def gen_task_stmt(self, app) -> str:
-5 """
-6 Generate the policy for placing tasks.
-7 Example generated code:
-8 Task * GPU; Task task1 CPU;
-9 """
-10 code = ""
-11 for task in app.tasks:
-12 proc = random.choice(["GPU","CPU"])
-[…]
-14 return code
-15
-16 def forward(self, app):
-17 task_stmt = self.gen_task_stmt(app)
-18 region_stmt = self.gen_region_stmt(app)
-19 layout_stmt = self.gen_layout_stmt(app)
-20 ...
-21 return task_stmt + region_stmt + ...
-(a) The decision procedures in generating a mapper.
+Table A1: Lines of Code (LoC) comparison between DSL and C++ mappers. The DSL achieves a 14×14$\times$14 × average reduction in LoC, making it a more suitable target for LLM code generation.
+A.3 Parallel Matrix Multiplication Algorithms
+2D Algorithms
+Cannon’s (Cannon, 1969) introduced a systolic communication pattern with tiled data partitioning for distributed matrix multiplication.
+PUMMA (Choi et al., 1994) and SUMMA (Van De Geijn & Watts, 1997) extended this approach by supporting non-square matrices and improving communication efficiency through pipelining.
+They are called 2D algorithms because they partition the matrices into 2D tiles and then map them onto the processor space.
+Non-2D Algorithms
+Johnson’s (Agarwal et al., 1995) introduced a 3D algorithm that partitions the input matrices into 3D tiles and uses additional memory per processor to reduce communication compared to 2D algorithms.
+Solomonik’s (Solomonik & Demmel, 2011) balances between 2D and 3D approaches by using extra memory to further minimize communication.
+COSMA (Kwasniewski et al., 2019) takes a different approach by optimizing the processor grid and parallelization strategy based on the input size and the machine size.
+A.4 Examples of Feedback Configurations
+We give examples for the raw execution output and enriched feedback in Table A2. The enhanced feedback includes explanations of errors and suggestions for mapper modifications.
+Case Raw Execution Output AutoGuide
+Explain Suggest
+case1 Compile Error: Syntax error, unexpected :, expecting { N/A There should be no colon : in function definition.
+case2 Compile Error: IndexTaskMap’s function undefined N/A Define the IndexTaskMap function first before using it.
+case3 Compile Error: mgpu not found N/A Include mgpu = Machine(GPU); in the generated code.
+case4 Execution Error: Assertion failed: stride does not match expected value. Memory layout is unexpected. Adjust the layout constraints or move tasks to different processor types.
+case5 Execution Error: DGEMM parameter number 8 had an illegal value Memory layout is unexpected. Adjust the layout constraint.
+case6 Execution Error: Slice processor index out of bound IndexTaskMap statements cause error.
+Ensure that the first index of mgpu ends with % mgpu.size[0], and the second element ends with % mgpu.size[1].
+case7 Execution Error: Assertion ‘event.exists()’ failed InstanceLimit statements cause error. Avoid generating InstanceLimit statements.
+case8 Performance Metric: Execution time is 0.03s. N/A Move more tasks to GPU to reduce execution time.
+case9 Performance Metric: Achieved throughput = 4877 GFLOPS N/A Try using different IndexTaskMap or SingleTaskMap statements to maximize throughput.
+Table A2: Raw execution output and AutoGuide (error explanations and adjustment suggestions) for different cases.
+A.5 Trace Agent Code
+Trace (Cheng et al., 2024) uses Python decorators like @bundle to annotate Python programs.
+It allows us to design an LLM code generation agent as if we were writing a Python program ourselves.
+We first set up an end-to-end runnable Python program that can generate a valid mapper program by randomly making decisions over the search space.
+We show the high-level structure of our Trace Mapper in Figure A3.
+Figure A2 shows how we incorporate the feedback from the execution to update the agent.
+At each optimization step, Trace will execute DSLMapperGenerator and collect the corresponding execution flow to build up a graph.
+Then it will make a call to an LLM to perform an update to any function that is decorated with @bundle(trainable=True).
+The DSLMapperGenerator is structured in the same way as providing a search space specified by the DSL, where an LLM optimizer can make decisions along the pre-designed axes.
+We note that this type of design is only enabled by recent developments like Trace and is much more challenging to do using older LLM-based frameworks.
 1policy = MapperAgent()
@@
 21 optimizer.step()
-(b) Learning policy using LLM-based optimizers.
-Figure 5: Example agent built with the interface offered by Trace.
-Figure 5(a) shows the mapper agent (each decision is defined as a function).
-Figure 5(b) shows how we incorporate the feedback from the execution to update the agent.
-The full agent code is shown in Figure A6.
-We conceptualize the mapper generation problem as an online optimization problem.
-Given a triplet $(\Theta,\omega,{\mathcal{T}})$, where $\Theta$ is the overall set of possible code that can be produced to generate a mapper; $\omega$ defines the optimization objective, which is to create a mapper that maximizes the throughput; and ${\mathcal{T}}:\Theta\rightarrow{\mathcal{F}},{\mathcal{G}}$ is the task that returns $f,g={\mathcal{T}}(\theta)$, where $f$ is the feedback from executing the generated mapper (the performance metric measured after using the generated mapper to serve the application code), and $g$ is the process graph of how the mapper is generated.
-In our particular setup, the performance of the mapper (throughput) is deterministic because system researchers have carefully controlled all possible randomness in the environment.
-If the parameter space is numerical, this online optimization problem can be solved through bandit algorithms (Lattimore & Szepesvári, 2020), reinforcement learning (Sutton & Barto, 2018), or Bayesian optimization (Snoek et al., 2012), but these methods are less effective when the parameter is text (mapper code).
-In this online optimization problem, we leverage the DSL to bring structure to the parameter space and make the online optimization more efficient.
-$\theta$ is the program code, $\omega$, and $f$ are expressed as text.
-We leverage an LLM as an optimizer to solve this optimization problem.
-LLM has been shown to have some ability to “optimize” by iteratively proposing better solutions when an explicit objective is described in the text.
-This phenomenon has been observed and leveraged by Yang et al.
-(2024a); Cheng et al.
-(2024); Yuksekgonul et al.
-(2024); Patel et al.
-(2024) to solve various optimization problems.
-We design our mapper agent to make various mapping decisions, from selecting processors for tasks to mapping index tasks onto the processor space.
-A high-level schema of the mapper agent is shown in Figure 4.
-Agent Design
-We express our mapper agent as a Python program (MapperAgent in Figure 5(a)).
-This program serves as the initial starting point of the optimization.
-In this mapper agent, we manually break down the task of generating a monolithic mapper into several independent code segments, allowing the agent to decide what code to generate for each segment.
-This strategy works because the design of our DSL removes unnecessary dependence between code segments.
-Our modularization approach mimics the least-to-most prompting (Zhou et al., 2022), and our efforts to use programs as prompts echo the program-as-thought approach (Chen et al., 2022).
-Optimization
-We illustrate our workflow design in Figure 5, with the full template in Figure A6.
-Using the framework Trace (Cheng et al., 2024), we only ask the LLM to generate the method decorated with @trace.bundle(trainable=True) at each iteration of optimization.
-During an optimization step (shown in Figure 5(b)), the LLM receives complete information about how the agent generated the previous mapper, as well as feedback from the system execution.
-The LLM optimizer needs feedback from the system to perform the optimization.
-The system feedback can be categorized into three types: (1) Compile Error, indicating a failure of the LLM to generate code capable of producing a valid DSL; (2) Execution Error, such as an application running out of GPU memory; and (3) Performance Metric, which refers to the application’s running time or throughput.
-The goal of the optimization process is to minimize running time, thereby increasing throughput.
-Examples are provided in the “System Feedback” column in Table 2.
-Mapper System Feedback Enhanced Feedback
-Explain Suggest
-mapper1 Compile Error: Syntax error, unexpected :, expecting { N/A There should be no colon : in function definition.
-mapper2 Execution Error: Assertion failed: stride does not match expected value. Memory layout is unexpected. Adjust the layout constraints or move tasks to different processor types.
-mapper3 Performance Metric: Execution time is 0.03s. N/A Move more tasks to GPU to reduce execution time.
-Table 2: System feedback and enhanced feedback for different mappers.
-Enhanced Feedback
-To accelerate the LLM’s search for high-performance mappers, we introduce enhanced feedback in addition to system feedback.
-Table 2 provides examples with enhanced feedback, with further examples available in Section A.7.
-There are two types of enhanced feedback: the first is to explain execution errors, providing more informative failure messages to guide LLMs in generating correct mapper; the second is to suggest modifications to mapping decisions to avoid errors or improve performance.
-Enhanced feedback is implemented via keyword matching, where system feedback triggers the corresponding explanations and suggestions for mapper improvements.
-We conduct an ablation study in Section 5.4 to demonstrate the effectiveness of enhanced feedback in our experiments.
-5 Evaluation
-Experiments are conducted on a GPU cluster where each node has two Intel 10-core E5-2640 v4 CPUs, 256G main memory, and four NVIDIA Tesla P100 GPUs. Regarding the LLM, we use gpt-4o-2024-08-06.
-5.1 Effectiveness of the DSL
-To evaluate the necessity and effectiveness of our DSL in mapper code generation, we compared the generation of mappers in C++ with DSL.
-We devised 10 mapping strategies, each described in natural language, to serve as test cases for code generation [^1] An example strategy is “aligning all data to 64 bytes in memory and utilizing Fortran ordering for multi-dimensional data.”.
-A complete list of these strategies is provided in Section A.9.
-It is important to note that the objective is not to optimize the performance of any specific application, but rather to assess whether the LLM can accurately generate a C++/DSL mapper based on a given strategy.
-We provided the same types of materials in the prompt to ensure a fair comparison.
-We provided documentation [^2] We used documentation covering all features of DSL, and pre-existing documentation for C++ mapper., example mapper programs, and starting code for both DSL and C++ interfaces.
-We measured the success rate of generating mappers that can pass compilation and pass test cases for each strategy.
-For C++ code generation, we enhanced the process by incorporating compiler feedback for additional attempts.
-The LLM received compiler error messages and iteratively refined the code up to 10 times.
-We use DSPy (Khattab et al., 2023) to build our generation pipeline.
-Code Generation Target Mapping Strategy Success Rate
-1 2 3 4 5 6 7 8 9 10
-C++ (single trial) 0%
-C++ (iterative refine) 0%
-DSL (single trial) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 80%
-Table 3: Compilation and Strategy Test.
-We show the success rate for code generation given 10 mapping strategies.
-Generating DSL code significantly outperforms generating C++ (either without or with compiler feedback).
-– fails to compile, ✗ compiles but fails the test, and $\checkmark$ passes the test.
-ResultTable 3 presents the performance of the LLM in mapper code generation in three settings: C++ single trial (without compiler feedback), C++ with compiler feedback, and DSL.
-We evaluate the LLM with 10 different mapping strategies.
-As observed, the DSL approach consistently outperforms the other settings, followed by C++ with compiler feedback, while C++ without feedback performs the worst.
-These results confirm that the LLM is poor at generating system-level C++ code and highlight the effectiveness and necessity of our DSL design in significantly enhancing the code generation capabilities for mapping.
-Failure Case Analysis The compilation errors in C++ mapper generation arise because LLMs are unable to manage the framework-specific contextual dependencies inherent to low-level system software programming.
-For instance, LLMs generate variable names or references that do not exist within the provided codebase.
-These issues are particularly challenging to resolve since we cannot supply the full header files to the LLMs due to token limits.
-In the cases where the code compiles but fails the test cases, the failures are exclusively in C++.
-The root cause lies in the complexity of implementing a C++ mapper, which requires a deep understanding and reasoning about the low-level mapping APIs.
-As outlined in the documentation, multiple APIs must work in concert to accomplish specific tasks, yet LLMs are unable to grasp this level of coordination.
-While compiler feedback can guide LLMs in avoiding trivial errors (e.g., such as generating non-existent variable names), it cannot bridge the gap in understanding the intricacies of low-level C++ mapping APIs.
-This limitation explains why all attempts at C++ code generation ultimately fail.
-In contrast, only two test cases fail in the DSL context on the single trial, both due to compilation errors stemming from incorrect usage of the DSL’s syntax.
-This is understandable given that the DSL is a newly designed language with limited training data available for the LLM.
-Thus, for performance optimization experiments, we do not attempt to have LLMs optimize over C++ mappers, as results sufficiently demonstrate that LLMs struggle to generate C++ code that meets the specifications of what a desired mapper should do, let alone explore a large search space of different mapping strategies.
-Figure 6: Normalized throughput for scientific applications comparing expert-written mappers, random mappers, best mappers found by Trace, and the average optimization trajectories of Trace and OPRO in 10 iterations across 5 runs.
-5.2 Accelerating Scientific Applications
-We use the following three scientific applications as our benchmarks.
-The circuit simulation benchmark (Bauer et al., 2012) models the behavior of an electrical circuit by simulating currents and voltages across interconnected nodes and wires.
-The stencil computation benchmark (Van der Wijngaart & Mattson, 2014) simulates a 2D grid where each point’s value is updated based on its neighbors using a stencil pattern.
-The Pennant benchmark (Ferenbaugh, 2015) models unstructured mesh, Lagrangian staggered-grid hydrodynamics, used for simulating compressible flow.
-In this experiment, we compared the performance of different mappers.
-All mappers are implemented using our DSL.
-Our goal is to determine whether LLMs can effectively explore the search space of mappers by generating high-performance DSL mappers.
-Expert-written mappers were manually developed and optimized by domain experts as part of the application development process.
-We re-implemented these expert-written C++ mappers using our DSL to establish a ground truth for comparison with our approach.
-Validation confirmed that the DSL-based mappers achieve performance equivalent to the original C++ mappers, providing a fair basis for evaluating the performance of our generated mappers.
-Randomly generated mappers are produced by our MapperAgent with 10 different random seeds and serve as a baseline for this experiment.
-We ran 10 randomly generated mappers and reported the average performance.
-LLM-generated mappers are implemented using Trace (Cheng et al., 2024), which employs LLMs as optimizers.
-We tested both the Trace and OPRO (Yang et al., ) search algorithms, running 10 iterations for each application.
-Due to the stochastic nature of LLM outputs, we repeated the optimization process 5 times and averaged the results.
-We also report the best mapper found by Trace across the 5 runs.
-The performance-critical mapping decisions for these applications are the processor type selection, memory placement of data, and the layout selection for data.
-Other decisions available to mappers do not change the performance of these applications much.
-The simplest application above (with the smallest search space) is Stencil, which contains 2 tasks and 12 data arguments.
-Each task and data argument has two choices of placement, together with 4 additional layout choices for each data argument, forming an optimization space of $2^{38}$.
-Results
-We use normalized throughput as the performance metric in Figure 6, where higher values indicate better performance.
-The throughput is normalized w.r.t the expert-written mappers.
-As shown in Figure 6, random mappers are consistently the least effective across all applications, highlighting the importance of mapping decisions on application performance.
-When comparing the optimization trajectories of Trace and OPRO, Trace performs similarly to OPRO, and is slightly better than OPRO on Pennant.
-All the best mappers found by Trace can at least match the performance of expert mappers.
-Figure 7: Normalized throughput for matrix-multiplication algorithms.
-We compare expert-written mappers, random mappers, best mappers found by Trace, and the average optimization trajectories of Trace and OPRO in 10 iterations across 5 runs.
-Interestingly, the best mapper identified in the Circuit benchmark outperforms the expert mapper by 1.34×1.34$\times$1.34 ×.
-Upon manual investigation, we observed that the key difference lies in memory placement: the best mapper allocates two data collections to the GPU FrameBuffer memory, whereas the expert mapper places these collections in GPU ZeroCopy memory.
-This strategy reduces task execution time, despite a slight increase in inter-GPU communication costs, ultimately leading to improved overall performance.
-For the Pennant benchmark, while there is a minor difference in data collection placement, the final performance results between the two mappers are nearly equivalent.
-For each application, the optimization process completes within 10 minutes, significantly reducing mapper development time from days to minutes.
-This substantial improvement highlights the efficiency of our LLM-enhanced DSL in quickly generating high-performance mappers, offering clear benefits for both developers and application performance.
-5.3 Accelerating Matrix Multiplications
-We follow the same experimental setup as described in Section 5.2, with the key difference being the focus on matrix multiplication algorithms in this subsection.
-Different from the applications in Section 5.2, the important mapping decision here is index mapping, which is the choice of how to launch concurrent tasks and distribute the tiles of matrices across multiple GPUs.
-In the DSL mapper, achieving a high-performance index mapping requires specifying a function that maps the iteration space (defined by the parallel for loop) to the processor space of the machine (discussed in Section 4.1 IndexTaskMap).
-Based on our estimates, each algorithm has approximately $10^{9}$ possible choices for index mapping.
-We show some of the mapping functions in Section A.5
-We target 6 different matrix multiplication algorithms: Cannon’s (Cannon, 1969), SUMMA (Van De Geijn & Watts, 1997), PUMMA (Choi et al., 1994), Johnson’s (Agarwal et al., 1995), Solomonik’s (Solomonik & Demmel, 2011), and COSMA (Kwasniewski et al., 2019).
-Each algorithm exhibits different performance characteristics and may be the preferred implementation of matrix multiply depending on the target machine and input size.
-Our goal is to explore and identify better mappings for these algorithms.
-We elaborate further on these algorithms in Section A.4.
-Results
-As shown in Figure 7, the algorithm self-specified expert mappers’ compute throughput is normalized to 1.0.
-Random mappers only yield 2%-40% of the throughput achieved by expert mappers, highlighting the importance of high-quality mapper design.
-The best mappers found by Trace consistently outperform the expert mappers, achieving speedups ranging from 1.09×1.09$\times$1.09 × to 1.31×1.31$\times$1.31 ×.
-For both PUMMA and Solomonik’s, the throughput of the best mapper discovered by Trace is significantly higher than the average optimization trajectory across 5 runs.
-This variability is due to the randomness of LLMs in our experiments, where even small changes in mappers can lead to notable performance differences.
-When comparing optimization trajectories of Trace and OPRO, Trace shows significant performance gains in SUMMA, PUMMA, Johnson’s, and COSMA, while achieving similar results on other benchmarks.
-Finally, we inspected and compared the mapping decisions of the expert mappers with those of the best mappers found through our search.
-The performance improvements are entirely attributed to more effective index mapping.
-Index mapping controls the partitioning and distribution of data, in this case, matrices, across the GPUs.
-The optimized mapping reduces inter-GPU communication and enhances data locality, leading to improved performance in parallel matrix multiplication algorithms.
-Figure 8: Comparison of three types of feedback design.
-System provides only the system execution information (performance numbers or error messages) to the optimizer.
-Explain provides additional explanations of execution errors.
-Suggest provides modification proposals to the mappers.
-All feedback is automatically generated.
-5.4 Ablation Study of Feedback
-Within the Trace framework, feedback is provided to the LLM optimizer to inform subsequent improvement on mappers.
-The quality of the feedback directly affects the success of the optimization process.
-We evaluate three types of feedback to assess how different feedback influences the performance of LLM-based optimizers during mapper search (described in Table 2): (1) system feedback (compile error, execution error, or performance metrics), (2) error explanations, and (3) suggestions for mapper adjustments.
-The first feedback message includes only system feedback (labeled System in Figure 8).
-Next, we evaluate an enhanced feedback message that includes both system feedback and error explanations (labeled System+Explain).
-Finally, we provide the full feedback message, including all three types (labeled System+Explain+Suggest), corresponding to the Trace results shown in Figure 6 and Figure 7.
-We evaluate on three benchmarks: the Circuit simulation benchmark and two matrix multiplication algorithms (COSMA and Cannon’s).
-As shown in Figure 8, across all three benchmarks, the full feedback message consistently achieves the highest throughput after 10 iterations, followed by the one without guidance.
-The system-only feedback performs the worst among the three.
-While the degree of impact from the feedback types varies across benchmarks, the results clearly demonstrate the critical role of high-quality feedback in guiding LLM optimizers to discover more efficient mappers.
-6 Conclusion
-In this paper, we addressed the challenges of automating mapper generation in task-based programming through the use of LLMs and a Domain-Specific Language (DSL).
-By designing a high-level DSL, we effectively simplified the complex task of generating low-level C++ mappers, enabling LLMs to handle mapper generation without deep knowledge of system intricacies.
-Additionally, we formulated the mapper generation task as a discrete optimization problem, leveraging LLM optimizers to explore a structured and constrained search space defined by the DSL.
-Our experimental results demonstrate the efficacy of this approach, with LLM-generated mappers achieving up to 1.34×1.34$\times$1.34 × speedup over expert-written mappers across nine benchmarks.
-For matrix multiplication algorithms, we observed a performance boost of up to 1.31×1.31$\times$1.31 ×.
-These findings show that our LLM-enhanced DSL significantly reduces development time from days to minutes, benefiting both human developers and the performance of parallel applications.
-Reproducibility Statement
-We provide comprehensive details regarding the setup of our benchmark, ensuring full reproducibility based on the provided information.
-We are planning to open-source the framework, as well as the code to reproduce all experiment results.
-We provide detailed documentation of the evaluation process, along with all the prompts that we use.
-We use gpt-4o-2024-08-06 with a max token of 16383 and default temperature.
-All software used in our evaluation (e.g., Legion) is open source.
-We acknowledge the potential difference in performance measurement across different platforms.
-Acknowledgments
-We thank Yuhui Zhang, Chenyang Yang, Genghan Zhang, Qizheng Zhang, Mohammad R.
-Fadiheh, Ed Chen, Mert Yuksekgonul, Jiaxin Ge, Tingyu Zhu, Simon Guo and Anne Ouyang for the feedback.
-We thank Ching-An Cheng and Ken Liu for their discussions and feedback.
-Appendix A Appendix
-A.1 DSL Grammar
-We show the DSL grammar in Figure A1.
-[…]
-[…]
-$\mathsf{IndexTaskMap}\;\mathbf{TaskName}\;\mathbf{var}\mathrel{\mid}$
-$\mathsf{SingleTaskMap}\;\mathbf{TaskName}\;\mathbf{var}\mathrel{\mid}$
-$\mathtt{FuncDef}\mathrel{\mid}\mathtt{TaskMapModifier}$
-[…]
-[…]
-[…]
-[…]
-$\mathsf{InstanceLimit}\;\mathbf{TaskName}\;\mathbf{int}\mathrel{\mid}\ldots$
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-$\mathtt{Expr}\;\mathsf{Op}\;\mathtt{Expr}\mathrel{\mid}(\mathtt{Expr})\mathrel{\mid}\mathtt{Expr}[\mathtt{Expr}]\mathrel{\mid}*\;\mathtt{Expr}$
-$\mathtt{Expr}\;\mathtt{?}\;\mathtt{Expr}:\mathtt{Expr}$
-Figure A1: Grammar of DSL. A DSL program is a list of statements, each of which controls one aspect of mapping.
-A.2 Processor Space Transformation for Index Mapping
-Transformation Semantics
-$m^{\prime}=m.\mathtt{split}(i,d)$ $m^{\prime}[a_{0},\ldots,a_{n+1}]\coloneq m[b_{0},\ldots,b_{n-1},b_{n}]$
-$b_{t}=\begin{cases}a_{t}\qquad&\quad t<i\\ a_{i}+a_{i+1}\cdot d\qquad&\quad t=i\\ a_{t+1}\qquad&\quad t>i\end{cases}$
-$m^{\prime}=m.\mathtt{merge}(p,q)$ $m^{\prime}[a_{0},\ldots,a_{n-1}]\coloneq m[b_{0},\ldots,b_{n-1},b_{n}]$
-$p<q$ $b_{t}=\begin{cases}a_{t}\qquad&\quad t<p\vee p<t<q\\ a_{p}\%m.size[p]\qquad&\quad t=p\\ a_{p}/m.size[p]\qquad&\quad t=q\\ a_{p-1}\qquad&\quad t>q\\ \end{cases}$
-$m^{\prime}=m.\mathtt{swap}(p,q)$ $m^{\prime}[a_{0},\ldots,a_{n-1}]\coloneq m[b_{0},\ldots,b_{n-1}]$
-$b_{t}=\begin{cases}a_{q}\qquad&\quad t=p\\ a_{p}\qquad&\quad t=q\\ a_{t}\qquad&\quad t\neq p\land t\neq q\\ \end{cases}$
-$m^{\prime}=m.\mathtt{slice}(i,low,high)$ $m^{\prime}[a_{0},\ldots,a_{n-1}]\coloneq m[b_{0},\ldots,b_{n-1}]$
-$0\leq low\leq high<m.size[i]$ $b_{t}=\begin{cases}a_{i}+low\qquad&\quad t=i\\ a_{t}\qquad&\quad t\neq i\\ \end{cases}$
-Figure A2: Semantics of processor space transformations expressed as mappings from the indices of the transformed processor space to the indices of the original processor space.
-A processor space $m$ can be initialized via m = Machine(GPU); as a 2D tuple where the first dimension represents the node index, and the second dimension represents the index of GPU within the node.
-To better express mapping between index spaces, we define a set of transformation primitives.
-A transformation primitive is a function of the processor space $m$ that returns a transformed processor space $m^{\prime}$, where $m$ and $m^{\prime}$ are related through the mapping shown in the right-hand side of Figure A2.
-Our transformations are inspired by widely used operations for changing the dimensionality of arrays in libraries such as NumPy, but the application to mapping is quite different.
-We now explain each transformation in detail.
-The split transformation takes two arguments: an integer $i$ that indicates the dimension to be split, and the splitting factor $d$.
-Suppose $m$ is a processor space of size $(8,8)$, then after executing $m^{\prime}=m.\mathtt{split}(0,2)$, $m^{\prime}$ will be a processor space of size $(2,4,8)$.
-An important property of split and all DSL transformations is that they are invertible.
-Thus, mappers can work with the transformed space $m^{\prime}$ but DSL can translate such uses back into the original processor space $m$ to identify which concrete processors to use.
-In this example, $m^{\prime}[j_{0},j_{1},j_{2}]=m[j_{0}+j_{1}\times 2,j_{2}]$.
-The merge transformation takes two dimensions of the original processor to be fused as its input.
-Suppose $m^{\prime}$ is a processor space of size $(2,4,8)$.
-After applying $m^{\prime\prime}=m^{\prime}.\mathtt{merge}(0,1)$, $m^{\prime\prime}$ will be a processor space of size $(8,8)$.
-The processor indexed by $m^{\prime\prime}[j_{0},j_{1}]$ corresponds to $m^{\prime}[j_{0}\%2,j_{0}/2,j_{1}]$.
-Transformations can be chained together.
-Suppose $m$ is a 2D processor space, and we start with $m^{\prime}=m.\mathtt{split}(0,d)$, followed by $m^{\prime\prime}=m^{\prime}.\mathtt{merge}(0,1)$.
-The final processor space $m^{\prime\prime}$ is a 2D processor space.
-Now we will derive the index transformation from $m^{\prime\prime}$ to $m$ by applying the transformation rules from merge transformation and split transformation: $m^{\prime\prime}[j_{0},j_{1}]=m^{\prime}[j_{0}\%d,j_{0}/d,j_{1}]=m[(j_{0}\%d)+(j_{0}/d)\times d,j_{1}]$.
-The expression $(j_{0}\%d)+(j_{0}/d)\times d$ can be simplified to $j_{0}$ because the division operator /// between two integers rounds to zero.
-Therefore, $m^{\prime\prime}[j_{0},j_{1}]=m[j_{0},j_{1}]$, showing that the split and merge transformation primitives are inverses of each other.
-The swap transformation primitive takes two parameters indicating the dimensions to be swapped and returns a processor space where the two indices of the chosen dimensions are flipped.
-The swap transformation is often combined with merge: The merge primitive linearizes two dimensions into one, but there is a choice whether to use row-major or column-major iteration order in the linearization.
-Users can change the iteration order for the merge by swapping the two dimensions.
-The slice transformation primitive takes three parameters, the dimension to slice, and the lower bound and upper bound of the dimension.
-The index mapping rule for the slice transformation is to add a constant shift in the chosen dimension.
-The slice transformation primitive can be useful if users want to map the iteration space to only part of the original processor space.
-If two iteration spaces can be executed concurrently in the program, users can map one iteration space to half of the processors and map the other iteration space to the other half of the processors.
-In this case, the slice transformation allows users to map an iteration space to the selected portion of the processor space.
-A.3 Common Index Mapping Functions
-We show the common index mapping functions in Figure A3.
-Figure A3: Common transformations and index mapping functions.
-The shaded subarea of the iteration space will be mapped to the shaded processor in the processor space.
-The transformation code can transform the original $(2,2)$ processor space into the desired processor space.
-The result processor space will be used for mapping in the user-defined function.
-A.4 Parallel Matrix Multiplication Algorithms
-2D Algorithms
-Cannon’s (Cannon, 1969) introduced a systolic communication pattern with tiled data partitioning for distributed matrix multiplication.
-PUMMA (Choi et al., 1994) and SUMMA (Van De Geijn & Watts, 1997) extended this approach by supporting non-square matrices and improving communication efficiency through pipelining.
-They are called 2D algorithms because they partition the matrices into 2D tiles and then map them onto the processor space.
-Non-2D Algorithms
-Johnson’s (Agarwal et al., 1995) introduced a 3D algorithm that partitions the input matrices into 3D tiles and uses additional memory per processor to reduce communication compared to 2D algorithms.
-Solomonik’s (Solomonik & Demmel, 2011) balances between 2D and 3D approaches by using extra memory to further minimize communication.
-COSMA (Kwasniewski et al., 2019) takes a different approach by optimizing the processor grid and parallelization strategy based on the input size and the machine size.
-A.5 Index mapping functions used by Matrix Multiplication Algorithms
-We show some index mapping functions used by matrix multiplication algorithms in Figure A4
-Figure A4: Example mapping functions used by the mappers of matrix multiplication algorithms.
-A.6 Explanation of Index Mapping for Solomonik’s Algorithm
-Figure A5: Mapper of the Solomonik’s algorithm on a 2-node machine with 4 GPUs per node.
-The 2D processor space is transformed via the split transformation primitive into a 6D space (visualized as two 3D spaces).
-Figure A5 shows a mapper for the Solomonik’s algorithm on a 2-node machine with 4 GPUs per node.
-The result distribution for the 3D iteration space is that each node will get half of the whole iteration space by partitioning along the x-axis, and the 4 GPUs per node will perform a 2D block distribution over the y-z plane.
-There is a dimension mismatch between the iteration space (3D) and the initial processor space (2D).
-To conduct the desired mapping required by the algorithm, we first apply the split transformation primitive four times (colored as red in the code).
-We apply the first (resp.
-last) two split transformations to make the node dimension (resp.
-GPU dimension) align with the 3D iteration space.
-We visualize the result 6D processor space as two 3D spaces.
-The first 3D space (representing the node dimension) is of size $(2,1,1)$ and the second 3D space (representing the GPU dimension) is of size $(1,2,2)$.
-A.7 Examples of Feedback Configurations
-We give examples for the system feedback and enhanced feedback in Table A1. The enhanced feedback includes explanations of errors and suggestions for mapper modifications.
-Mapper System Feedback Enhanced Feedback
-Explain Suggest
-mapper1 Compile Error: Syntax error, unexpected :, expecting { N/A There should be no colon : in function definition.
-mapper2 Compile Error: IndexTaskMap’s function undefined N/A Define the IndexTaskMap function first before using it.
-mapper3 Compile Error: mgpu not found N/A Include mgpu = Machine(GPU); in the generated code.
-mapper4 Execution Error: Assertion failed: stride does not match expected value. Memory layout is unexpected. Adjust the layout constraints or move tasks to different processor types.
-mapper5 Execution Error: DGEMM parameter number 8 had an illegal value Memory layout is unexpected. Adjust the layout constraint.
-mapper6 Execution Error: Slice processor index out of bound IndexTaskMap statements cause error.
-Ensure that the first index of mgpu ends with % mgpu.size[0], and the second element ends with % mgpu.size[1].
-mapper7 Execution Error: Assertion ‘event.exists()’ failed InstanceLimit statements cause error. Avoid generating InstanceLimit statements.
-mapper8 Performance Metric: Execution time is 0.03s. N/A Move more tasks to GPU to reduce execution time.
-mapper9 Performance Metric: Achieved throughput = 4877 GFLOPS N/A Try using different IndexTaskMap or SingleTaskMap statements to maximize throughput.
-Table A1: System feedback and enhanced feedback (error explanations and adjustment suggestions) for different mappers.
-A.8 Trace Agent Code
-Trace (Cheng et al., 2024) uses Python decorators like @bundle to annotate Python programs.
-It allows us to design an LLM code generation agent as if we were writing a Python program ourselves.
-We first set up an end-to-end runnable Python program that can generate a valid mapper program by randomly making decisions over the search space.
-We show the high-level structure of our Trace Mapper in Figure A6.
-At each optimization step, Trace will execute DSLMapperGenerator and collect the corresponding execution flow to build up a graph.
-Then it will make a call to an LLM to perform an update to any function that is decorated with @bundle(trainable=True).
-The DSLMapperGenerator is structured in the same way as providing a search space specified by the DSL, where an LLM optimizer can make decisions along the pre-designed axes.
-We note that this type of design is only enabled by recent developments like Trace and is much more challenging to do using older LLM-based frameworks.
+Figure A2: We show how we use Trace to incorporate the feedback from the execution to update the agent, with a Pytorch-like syntax.
 1import opto.trace as trace
@@
 For each task, we produce a mapper from this starting agent and then ask LLMs to “optimize” this agent (by changing functions that are trainable) to produce mappers that are optimal for the particular task.
-A.9 Mapping Strategies
+A.6 Mapping Strategies
 Strategy 1: Map the tasks of calculate_new_currents, distribute_charge, update_voltages onto GPUs in this way: linearize the 2D GPU processor space into 1D, then perform 1D block mapping from launch domain to the linearized 1D processor space.
@@
 11Layout * * * Align==64 F_order;
-Strategy 6 Place the task calculate_new_currents onto CPU
+Strategy 6: Place the task calculate_new_currents onto CPU
 1Task * GPU,CPU; # for any task, run on GPU if supported
@@
 18IndexTaskMap calculate_new_currents,distribute_charge,update_voltages cyclic1d;
-A.10 Generated Mapper Examples
+A.7 Generated Mapper Examples
 Here we provide examples of generated mappers for a subset of problems.

### v4

@@
-Improving Parallel Program Performance with LLM Optimizers via Agent-System Interface
+Improving Parallel Program Performance with LLM Optimizers via Agent-System Interfaces
 Abstract
@@
 We introduce a framework that automates mapper development with generative optimization, leveraging richer feedback beyond scalar performance metrics.
-Our approach features the Agent-System Interface, which includes a Domain-Specific Language (DSL) to abstract away low-level complexity of system code and define a structured search space, as well as AutoGuide, a mechanism that interprets raw execution output into actionable feedback.
+Our approach features the Agent-System Interface, which includes a Domain-Specific Language (DSL) to abstract away the low-level complexity of system code and define a structured search space, as well as AutoGuide, a mechanism that interprets raw execution output into actionable feedback.
 Unlike traditional reinforcement learning methods such as OpenTuner, which rely solely on scalar feedback, our method finds superior mappers in far fewer iterations.
@@
 Unlike traditional methods such as reinforcement learning (Ansel et al., 2014), which rely solely on scalar rewards, generative optimization can utilize richer forms of feedback, such as error explanations and actionable suggestions expressed in natural language.
-This agentic optimization workflow has previously proven to be effective across various domains (Nie et al., 2024; Cheng et al., 2024; Yang et al., ; Khattab et al., 2023; Yuksekgonul et al., 2024).
+This agentic optimization workflow has previously proven to be effective across various domains (Nie et al., 2024; Cheng et al., 2024; Yang et al., 2023; Khattab et al., 2023; Yuksekgonul et al., 2024).
 Our work is the first to apply such technique to the domain of system optimization.
@@
 3 Problem Definition
-Motivation and Challenges
-The concrete problem we address is the automated generation of high-performance mappers for the Legion parallel programming framework (Bauer et al., 2012).
-Mappers dictate task scheduling and data placement.
-A well-designed mapper can achieve orders-of-magnitude speedup over naive strategies.
-However, automating mapper generation is challenging due to two key factors.
-First, the complexity of low-level system code.
-Implementing a mapper requires writing hundreds of lines of intricate C++ code, demanding expertise in system internals.
-Second, the vast search space of mapping strategies.
-The search space grows exponentially with the number of tasks and arguments.
-Search Space and Performance Impact
-As illustrated in Figure A1, the search space of mappers involves multiple decisions, each influencing performance.
-The first key aspect is processor selection, which determines whether a task runs on GPUs, CPUs, or the OpenMP runtime.
-This choice depends on factors such as task size, GPU memory capacity, and kernel launch overhead.
-For instance, small tasks may prefer CPUs due to the overhead of launching GPU kernels, while tasks with large memory footprints may run on CPUs when GPU memory is insufficient.
-Another crucial dimension is memory placement, which dictates where data is stored.
-A mapper must decide whether to place data in the GPU’s FrameBuffer for fast access, ZeroCopy memory for CPU-GPU sharing, or CPU system memory for more available storage.
-Each option presents trade-offs between access speed, memory usage, and data transfer overhead.
-Additionally, memory layout further expands the search space, with decisions on Struct of Arrays (SOA) vs.
-Array of Structures (AOS), data ordering (Fortran-order vs.
-C-order), and alignment constraints (e.g., 128-byte alignment) significantly affecting cache efficiency and performance.
-Finally, an important idiom in high-performance computing is launching tasks over partitioned data.
-Index mapping determines how data partitions and task executions are distributed across multiple processors.
-For consistency, we can represent data partitioning as a tensor of data partitions, the machine as a tensor of processors, and tasks operating on the partitioned data as a tensor of tasks.
-The way data and task indices are mapped to processor indices affects inter-processor communication, a key factor in performance (Unger et al., 2022; Zheng et al., 2022a).
-4 Our Approach: Agent-System Interface
 1# Map task0 to GPU.
@@
 The highlighted boxes illustrate how the same functionality, which requires extensive C++ system code, can be expressed concisely in just a few lines in DSL.
+Motivation and Challenges
+The concrete problem we address is the automated generation of high-performance mappers for the Legion parallel programming framework (Bauer et al., 2012).
+Mappers dictate task scheduling and data placement.
+A well-designed mapper can achieve orders-of-magnitude speedup over naive strategies.
+However, automating mapper generation is challenging due to two key factors.
+First, the complexity of low-level system code.
+Implementing a mapper requires writing hundreds of lines of intricate C++ code, demanding expertise in system internals.
+Second, the vast search space of mapping strategies.
+The search space grows exponentially with the number of tasks and arguments.
+Search Space and Performance Impact
+As illustrated in Figure A1, the search space of mappers involves multiple decisions, each influencing performance.
+The first key aspect is processor selection, which determines whether a task runs on GPUs, CPUs, or the OpenMP runtime.
+This choice depends on factors such as task size, GPU memory capacity, and kernel launch overhead.
+For instance, small tasks may prefer CPUs due to the overhead of launching GPU kernels, while tasks with large memory footprints may run on CPUs when GPU memory is insufficient.
+Another crucial dimension is memory placement, which dictates where data is stored.
+A mapper must decide whether to place data in the GPU’s FrameBuffer for fast access, ZeroCopy memory for CPU-GPU sharing, or CPU system memory for more available storage.
+Each option presents trade-offs between access speed, memory usage, and data transfer overhead.
+Additionally, memory layout further expands the search space, with decisions on Struct of Arrays (SOA) vs.
+Array of Structures (AOS), data ordering (Fortran-order vs.
+C-order), and alignment constraints (e.g., 128-byte alignment) significantly affecting cache efficiency and performance.
+Finally, an important idiom in high-performance computing is launching tasks over partitioned data.
+Index mapping determines how data partitions and task executions are distributed across multiple processors.
+For consistency, we can represent data partitioning as a tensor of data partitions, the machine as a tensor of processors, and tasks operating on the partitioned data as a tensor of tasks.
+The way data and task indices are mapped to processor indices affects inter-processor communication, a key factor in performance (Unger et al., 2022; Zheng et al., 2022a).
+4 Our Approach: Agent-System Interface
 4.1 Domain-Specific Language Design
@@
 To implement it, we develop a compiler that translates DSL into the low-level C++ APIs.
+Figure 3: Agent Optimization Process.
+The mapper agent takes server specifications and application-specific information as input, generates mapper code, and executes it alongside the application on the server.
+Raw execution feedback is enriched using the AutoGuide mechanism, and the mapper is iteratively refined by an LLM optimizer to improve performance.
+Case Raw Execution Output AutoGuide
+Explain Suggest
+Case 1 Execution Error: Assertion failed: stride does not match expected value. Memory layout is unexpected. Adjust the layout constraints or move tasks to different processor types.
+Case 2 Performance Metric: Execution time is 0.03s. N/A Move more tasks to GPU to reduce execution time.
+Table 1: AutoGuide Feedback Mechanism.
+The AutoGuide mechanism interprets raw execution output from the runtime system, providing more informative error explanations and suggestions for mapper modifications.
+It is implemented via keyword matching.
+Additional examples are shown in Table A3.
 As illustrated in Figure 2, the complexity of DSL code is significantly lower than that of C++.
@@
 In contrast, Figure 2(b) shows a snippet from a C++ mapper, emphasizing the intricacy of low-level implementation details.
-According to Table A1, using the DSL results in an average lines of code reduction of 14×14$\times$14 ×.
+Across the benchmarks, using the DSL results in an average lines of code reduction of 14×14$\times$14 ×.
 This substantial reduction makes DSL a more suitable target for LLM code generation, as it abstracts away the complexities inherent in low-level systems.
@@
 This decision applies to each task group launched by parallel for loops.
+Our DSL is designed to express a wide range of high-performance mapping strategies, including all of the most important decisions.
+While there may be cases where certain optimizations are not directly expressible, we have not encountered any.
+Despite being more constrained than general-purpose C++, the DSL has been proven to be effective: all mappers discovered by our agent that outperform expert-written C++ implementations are expressible within the current DSL.
 4.2 Generative Optimization via AutoGuide
-Figure 3: Agent Optimization Process.
-The mapper agent takes server specifications and application-specific information as input, generates mapper code, and executes it alongside the application on the server.
-Raw execution feedback is enriched using the AutoGuide mechanism and iteratively refined by an LLM optimizer to improve performance.
-Case Raw Execution Output AutoGuide
-Explain Suggest
-Case 1 Execution Error: Assertion failed: stride does not match expected value. Memory layout is unexpected. Adjust the layout constraints or move tasks to different processor types.
-Case 2 Performance Metric: Execution time is 0.03s. N/A Move more tasks to GPU to reduce execution time.
-Table 1: AutoGuide Feedback Mechanism.
-The AutoGuide mechanism interprets raw execution output from the runtime system, providing more informative error explanations and suggestions for mapper modifications.
-It is implemented via keyword matching.
-Additional examples are shown in Table A2.
 We formulate mapper generation as an online optimization problem.
@@
 We present the optimization process in Figure 3.
-The agent takes two inputs: server specifications (e.g., CPU/GPU counts) and application information (e.g., task lists, data arguments).
-It generates mapper code, which is executed alongside the application code on the server.
+The agent takes two inputs: server specifications and application metadata.
+Server specifications detail the hardware configuration, including the number of CPUs and GPUs per node, as well as the total node count.
+Application metadata provides information on task names and the associated data arguments accessed by each task.
+These inputs define the structured search space explored by the agent during optimization.
+The agent, using the given inputs, generates mapper code that is executed alongside the application code on the server.
 Raw execution feedback from the runtime is augmented with the AutoGuide mechanism and fed back to the LLM, iteratively refining the agent for improved mapper code generation.
+Figure 4: Performance Comparison.
+Normalized throughput for 9 benchmarks, comparing expert mappers, random mappers, the average optimization trajectories of Trace, OPRO, and OpenTuner in 10 iterations across 5 runs, and the best mappers found by Trace.
+Figure 5: Comparison of Trace (generative optimizer) and OpenTuner (traditional RL) over 1K iterations (averaged across all 9 benchmarks)
 Coding Agent
@@
 5.1 Speedup of Application Performance
-Figure 4: Performance Comparison.
-Normalized throughput for 9 benchmarks, comparing expert mappers, random mappers, the average optimization trajectories of Trace, OPRO, and OpenTuner in 10 iterations across 5 runs, and the best mappers found by Trace.
-Figure 5: Comparison of Trace (generative optimizer) and OpenTuner (traditional RL) over 1K iterations (averaged across all 9 benchmarks)
-We evaluate our approach using 9 benchmarks.
-Circuit (Bauer et al., 2012) is a simulation benchmark that models electrical circuit behavior by simulating currents and voltages across interconnected nodes and wires.
-Stencil (Van der Wijngaart & Mattson, 2014) simulates a 2D grid where each point’s value is updated based on a stencil pattern determined by its neighbors.
-Pennant (Ferenbaugh, 2015) models unstructured mesh Lagrangian staggered-grid hydrodynamics, commonly used for simulating compressible flow.
-The remaining six benchmarks – Cannon’s (Cannon, 1969), SUMMA (Van De Geijn & Watts, 1997), PUMMA (Choi et al., 1994), Johnson’s (Agarwal et al., 1995), Solomonik’s (Solomonik & Demmel, 2011), and COSMA (Kwasniewski et al., 2019) – are widely studied parallel matrix multiplication algorithms, which we discuss in more detail in Section A.3.
+Benchmarks
+Our evaluation utilizes a suite of 9 benchmarks, including 3 scientific computing workloads and 6 well-known matrix multiplication algorithms.
+Circuit is a simulation benchmark that models electrical circuit behavior by simulating currents and voltages across interconnected nodes and wires (Bauer et al., 2012).
+Stencil simulates a 2D grid where each point’s value is updated based on a stencil pattern determined by its neighbors (Van der Wijngaart & Mattson, 2014).
+Pennant models unstructured mesh Lagrangian staggered-grid hydrodynamics, commonly used for simulating compressible flow (Ferenbaugh, 2015).
+The remaining six benchmarks – Cannon’s, SUMMA, PUMMA, Johnson’s, Solomonik’s, and COSMA – are well-known parallel matrix multiplication algorithms (Cannon, 1969; Van De Geijn & Watts, 1997; Choi et al., 1994; Agarwal et al., 1995; Solomonik & Demmel, 2011; Kwasniewski et al., 2019).
+Parallel matrix multiplication remains an active research topic due to its central role in high-performance computing and scientific simulations (Yadav et al., 2022).
+Furthermore, improving matrix multiplication performance has a broad impact, as it accelerates numerous downstream machine learning workloads (Jangda et al., 2022; Zheng et al., 2025).
+We discuss these matrix multiplication algorithms in more detail in Section A.3.
+This benchmark suite provides both depth with its representative matrix multiplication algorithms and variety with its range of scientific computing workloads.
 In this experiment, we evaluate the performance of the mappers with the following baselines.
@@
 Agent-Optimized Mappers.
-Using Trace (Cheng et al., 2024), we evaluated the Trace and OPRO (Yang et al., ) search algorithms, running 10 iterations per application.
+Using Trace (Cheng et al., 2024), we evaluated the Trace and OPRO (Yang et al., 2023) search algorithms, running 10 iterations per application.
 To account for stochastic output, we repeated the process 5 times and report the average.
@@
 Results
-We use normalized throughput as the performance metric in Figure 4, where higher values indicate better performance.
-The throughput is normalized relative to the expert-written mappers.
+In Figure 4, we use normalized throughput as our performance metric, where higher values indicate better performance.
+The throughput is normalized relative to the expert-written mappers, providing a clear baseline for comparison.
+Our focus is on measuring end-to-end performance, which includes both the correctness and efficiency of the generated mappers.
+If the generated code has any syntax or runtime issues, its throughput is recorded as 0.
+We report the best mappers found by Trace, and the average optimization trajectories of Trace, OPRO and OpenTuner over 10 iterations across 5 runs.
 All the best mappers found by Trace can match or surpass the expert-written mappers, underscoring the effectiveness of agent-based generative optimizer.
-Random mappers consistently exhibit low performance across all applications, emphasizing the critical role of mapping decisions.
-When comparing optimization trajectories, Trace performs similarly to OPRO, and significantly outperforms OpenTuner.
-To further compare the agent-based optimizer with traditional reinforcement learning, we extended OpenTuner’s optimization iterations from 10 to 1000, as shown in Figure 5, where the x-axis is the log-scale of iterations and the y-axis represents relative throughput (averaged across all 9 benchmarks).
-Notably, Trace achieves a 3.8×3.8$\times$3.8 × speedup over OpenTuner even when OpenTuner is run for 1000 iterations.
-When both are limited to 10 iterations, Trace outperforms OpenTuner by $\times1$, demonstrating its ability to quickly identify high-performance mappings.
-This highlights the superiority of Trace (generative optimizer) over OpenTuner (traditional reinforcement learning).
-Moreover, Trace completes the entire optimization process in just 10 minutes per application, reducing mapper development time from days to minutes.
-Case Analysis
-The largest performance gain achieved by Trace over the expert mapper is observed in Circuit, with a speedup of 1.34×1.34$\times$1.34 ×.
-This improvement is primarily due to memory placement: the best mapper allocates two data collections to GPU FrameBuffer memory, while the expert mapper places them in GPU ZeroCopy memory.
-Despite a slight increase in inter-GPU communication costs, Trace reduces task execution time due to faster memory access, resulting in higher overall performance.
-For matrix-multiplication algorithms, the greatest speedup is seen in COSMA, with Trace achieving a 1.31×1.31$\times$1.31 × speedup over the expert mapper.
-This is attributed to Trace’s more efficient index mapping functions, which reduce inter-GPU communication by better distributing partitioned submatrices across GPUs.
-For additional context, examples of Trace mappers are presented in Section A.7.
+In our context, reporting the best-performing mapper is appropriate.
+Mapper optimization is an offline process, and in practice, it is standard to run the optimizer multiple times and deploy the best result.
+Once identified, the mapper can be reused across repeated executions on the same application, input, and hardware, incurring no further search cost.
 Code Generation Target Mapping Strategy Success Rate
@@
 Symbols indicate results: – fails to compile, ✗ compiles but fails the test, and $\checkmark$ passes the test.
+Random mappers consistently exhibit low performance across all applications, emphasizing the critical role of mapping decisions.
+For each application, we generate 10 random mappers by sampling from the full DSL-defined search space, totaling 90 mappers across 9 applications.
+Among them, 74 (82.2%) raise runtime errors due to invalid mapping decisions.
+The runtime system enforces correctness by rejecting such mappers during execution, resulting in a throughput of 0.
+When comparing optimization trajectories, Trace performs similarly to OPRO, and significantly outperforms OpenTuner.
+To further compare the agent-based optimizer with traditional reinforcement learning, we extended OpenTuner’s optimization iterations from 10 to 1000, as shown in Figure 5, where the x-axis is the log-scale of iterations and the y-axis represents relative throughput (averaged across all 9 benchmarks).
+Notably, Trace achieves a 3.8×3.8$\times$3.8 × speedup over OpenTuner even when OpenTuner is run for 1000 iterations.
+When both are limited to 10 iterations, Trace outperforms OpenTuner by $\times1$, demonstrating its ability to quickly identify high-performance mappings.
+This highlights the superiority of Trace (generative optimizer) over OpenTuner (traditional reinforcement learning).
+Moreover, Trace completes the entire optimization process in just 10 minutes per application, reducing mapper development time from days to minutes.
+To offer a more comprehensive view of performance variations, we present additional statistics, including the mean, standard deviation, worst, median, and best normalized throughput for both our method and OpenTuner.
+These statistics are derived from five runs for each benchmark, as detailed in Section A.4.
+Case Analysis
+The largest performance gain achieved by Trace over the expert mapper is observed in Circuit, with a speedup of 1.34×1.34$\times$1.34 ×.
+This improvement is primarily due to memory placement: the best mapper allocates two data collections to GPU FrameBuffer memory, while the expert mapper places them in GPU ZeroCopy memory.
+Despite a slight increase in inter-GPU communication costs, Trace reduces task execution time due to faster memory access, resulting in higher overall performance.
+For matrix-multiplication algorithms, the greatest speedup is seen in COSMA, with Trace achieving a 1.31×1.31$\times$1.31 × speedup over the expert mapper.
+This is attributed to Trace’s more efficient index mapping functions, which reduce inter-GPU communication by better distributing partitioned submatrices across GPUs.
+For additional context, examples of Trace mappers are presented in Section A.8.
+5.2 Ablation Study of DSL for Code Generation
+In Section 5.1, we demonstrate the overall effectiveness of our approach.
+Here, we conduct an ablation study on the DSL, the core of the Agent-System Interface.
+Since successful generation is the foundation of optimization, this subsection focuses on how well the DSL helps LLMs generate syntactically and semantically correct mappers compared to C++, rather than directly optimizing performance.
 Figure 6: Comparison of different feedback designs.
@@
 All feedback is automatically generated.
-5.2 Ablation Study of DSL for Code Generation
-In Section 5.1, we demonstrate the overall effectiveness of our approach.
-Here, we conduct an ablation study on the DSL, the core of the Agent-System Interface.
-Since successful generation is the foundation of optimization, this subsection focuses on how well the DSL helps LLMs generate correct mappers compared to C++, rather than directly optimizing performance.
 Experiment Setup
@@
 This demonstrates the effectiveness of DSL’s design in abstracting system complexity and providing a high-level interface that enables LLMs to tackle complex system challenges in code generation.
-Incorporating iterative refinement with compiler feedback further improves success rates, resolving three compilation errors in C++ and two in the DSL.
+Incorporating iterative refinement with compiler feedback further improves success rates, resolving four compilation errors in C++ and two in the DSL.
 However, the gap between DSL mappers and C++ mappers remains substantial.
@@
 Second, the DSL reduces the amount of code.
-As shown in Table A1, LLMs achieve an average reduction of 14×14$\times$14 × in lines of code, simplifying code generation.
+As discussed before, LLMs achieve an average reduction of 14×14$\times$14 × in lines of code, simplifying code generation.
 These results underscore the importance of a high-level agent-system interface.
@@
 The 0-shot and 5-shot results perform the worst, underscoring the importance of feedback-based iterative refinement.
-This highlights the value of an agentic workflow, showing that performance improvements are not solely driven by prompting the LLM but are a direct result of the iterative refinement inherent in the workflow design.
+This highlights the value of an agentic workflow, showing that performance improvements are not solely driven by prompting the LLM but are a direct result of the iterative refinement in the workflow design.
 6 Conclusion
@@
 Acknowledgements
-We thank Yuhui Zhang, Chenyang Yang, Genghan Zhang, Qizheng Zhang, Mohammad R.
-Fadiheh, Ed Chen, Mert Yuksekgonul, Jiaxin Ge, Xinyan Hu, Tingyu Zhu, Simon Guo, Anne Ouyang, and Adhish Ramkumar for the feedback.
+This work was supported in part by a Google Research Award.
+We thank Yuhui Zhang, Genghan Zhang, Qizheng Zhang, Mohammad R.
+Fadiheh, Ed Chen, Mert Yuksekgonul, Chenyang Yang, Jiaxin Ge, Simon Guo, and Anne Ouyang for the feedback.
 We thank Ching-An Cheng and Ken Liu for their discussions and feedback.
 Impact Statement
-This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+This paper presents work whose goal is to advance the field of Machine Learning.
+There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
 Appendix A Appendix
@@
 Figure A1: Mappers decide the placement of each task in the task graph to processors, the placement of data to memory, and how the iteration space of data is partitioned and mapped to different processors.
-A.2 Tables for Lines of Code Comparison
-We show the lines of code comparison between DSL and C++ mappers in Table A1.
-Application 1 2 3 4 5 6 7 8 9 Avg.
-LoC in C++ 347 306 379 447 437 430 428 433 448 406
-LoC in DSL 16 14 16 38 38 38 33 38 32 29
-LoC Reduction $\times2$ $\times2$ 24×24$\times$24 × 12×12$\times$12 × 12×12$\times$12 × $\times1$ 13×13$\times$13 × $\times1$ 14×14$\times$14 × 14×14$\times$14 ×
-Table A1: Lines of Code (LoC) comparison between DSL and C++ mappers. The DSL achieves a 14×14$\times$14 × average reduction in LoC, making it a more suitable target for LLM code generation.
+A.2 DSL Grammar
+Terminals: TaskName, RegionName, var, int
+Grammar Rules:
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 A.3 Parallel Matrix Multiplication Algorithms
@@
 COSMA (Kwasniewski et al., 2019) takes a different approach by optimizing the processor grid and parallelization strategy based on the input size and the machine size.
-A.4 Examples of Feedback Configurations
+A.4 Additional Performance Statistics
+In our setting, reporting the best result across multiple runs is appropriate, as the best mapper is the one that is desired by the user.
+Mapper search is an offline optimization process, and it is feasible to run the optimizer multiple times to select the highest-performing mapper.
+Once identified, this mapper can be reused without incurring additional search cost, as the deployment scenario (application, input, and hardware) remains fixed.
+That said, additional statistics on performance varations can provide a more complete picture.
+Here we include the mean, standard deviation, worst, median, and best normalized throughput across five runs for each benchmark of our approach Trace and OpenTuner.
+Benchmark Mean Std Dev Worst Median Best
+Circuit 1.33$\times$ 0.01 1.31$\times$ 1.33$\times$ 1.34$\times$
+Stencil 1.01$\times$ 0.01 1.00$\times$ 1.01$\times$ 1.02$\times$
+Pennant 1.03$\times$ 0.02 1.00$\times$ 1.03$\times$ 1.04$\times$
+Cannon 1.09$\times$ 0.00 1.08$\times$ 1.09$\times$ 1.09$\times$
+SUMMA 0.86$\times$ 0.48 0.00$\times$ 1.07$\times$ 1.09$\times$
+PUMMA 0.57$\times$ 0.55 0.00$\times$ 0.66$\times$ 1.09$\times$
+Johnson 0.98$\times$ 0.17 0.68$\times$ 1.06$\times$ 1.07$\times$
+Solomonik 0.52$\times$ 0.41 0.00$\times$ 0.61$\times$ 1.09$\times$
+COSMA 1.25$\times$ 0.03 1.23$\times$ 1.23$\times$ 1.31$\times$
+Table A1: Normalized throughput of our framework Trace.
+Benchmark Mean Std Dev Worst Median Best
+Circuit 0.97$\times$ 0.16 0.81$\times$ 0.99$\times$ 1.20$\times$
+Stencil 0.00$\times$ 0.00 0.00$\times$ 0.00$\times$ 0.00$\times$
+Pennant 0.00$\times$ 0.00 0.00$\times$ 0.00$\times$ 0.00$\times$
+Cannon 0.00$\times$ 0.00 0.00$\times$ 0.00$\times$ 0.00$\times$
+SUMMA 0.00$\times$ 0.00 0.00$\times$ 0.00$\times$ 0.00$\times$
+PUMMA 0.00$\times$ 0.00 0.00$\times$ 0.00$\times$ 0.00$\times$
+Johnson 0.00$\times$ 0.00 0.00$\times$ 0.00$\times$ 0.00$\times$
+Solomonik 0.00$\times$ 0.00 0.00$\times$ 0.00$\times$ 0.00$\times$
+COSMA 0.00$\times$ 0.00 0.00$\times$ 0.00$\times$ 0.00$\times$
+Table A2: Normalized throughput of OpenTuner.
+Our method achieves relatively stable performance across most benchmarks.
+The higher variance and occasional 0.00× worst-case throughput observed in SUMMA, PUMMA, and Solomonik are due to invalid mapper configurations in the search space (e.g., violating cuBLAS layout constraints).
+The runtime enforces correctness by rejecting such configurations during execution.
+While the generative optimizer typically learns to avoid these cases through the AutoGuide mechanism, occasional failures within the 10-iteration budget are still possible.
+In practice, such failures can be mitigated by repeating the optimization and selecting the best-performing mapper.
+In contrast, OpenTuner, despite running the same number of iterations, fails to generate valid mappers for 8 out of 9 benchmarks.
+This highlights the difficulty of exploring the search space using traditional reinforcement learning methods.
+A.5 Examples of Feedback Configurations
 We give examples for the raw execution output and enriched feedback in Table A3. The enhanced feedback includes explanations of errors and suggestions for mapper modifications.
@@
 Table A3: Raw execution output and AutoGuide (error explanations and adjustment suggestions) for different cases.
-A.5 Trace Agent Code
+A.6 Trace Agent Code
 Trace (Cheng et al., 2024) uses Python decorators like @bundle to annotate Python programs.
@@
 For each task, we produce a mapper from this starting agent and then ask LLMs to “optimize” this agent (by changing functions that are trainable) to produce mappers that are optimal for the particular task.
-A.6 Mapping Strategies
+A.7 Mapping Strategies
 Strategy 1: Map the tasks of calculate_new_currents, distribute_charge, update_voltages onto GPUs in this way: linearize the 2D GPU processor space into 1D, then perform 1D block mapping from launch domain to the linearized 1D processor space.
@@
 18IndexTaskMap calculate_new_currents,distribute_charge,update_voltages cyclic1d;
-A.7 Generated Mapper Examples
+A.8 Generated Mapper Examples
 Here we provide examples of generated mappers for a subset of problems.
@@
 6
-7def block1d(Task task) {
-8 ip = task.ipoint;
-9 return mgpu[ip[0] % mgpu.size[0], ip[0] % mgpu.size[1]];
-10}
-11
-12IndexTaskMap task_1 block1d;
-13
-14def cyclic1d(Task task) {
-15 ip = task.ipoint;
-16 linearize = ip[0] * 2 + ip[1];
-17 return mgpu[ip[0] % mgpu.size[0], linearize % mgpu.size[1]];
-18}
-19
-20IndexTaskMap task_1 cyclic1d;
-21
-22def cyclic2d(Task task) {
-23 ip = task.ipoint;
-24 linearize = ip[0] + ip[1] * 2;
-25 return mgpu[ip[0] % mgpu.size[0], linearize % mgpu.size[1]];
-26}
-27
-28IndexTaskMap task_1 cyclic2d;
-29
-30def linearize3D(Task task) {
+7def linearize3D(Task task) {
 8 ip = task.ipoint;
@@
 10 return mgpu[linearize % mgpu.size[0], linearize % mgpu.size[1]];
-34}
-35
-36IndexTaskMap task_1 linearize3D;
-37
-38def linearize2D(Task task) {
+11}
+12
+13IndexTaskMap task_1 linearize3D;
+14
+15def linearize2D(Task task) {
 16 ip = task.ipoint;
@@
 18 return mgpu[linearize % mgpu.size[0], linearize % mgpu.size[1]];
-42}
-43
-44IndexTaskMap task_1 linearize2D;
-45IndexTaskMap task_2 block1d;
-46IndexTaskMap task_2 cyclic1d;
-47IndexTaskMap task_2 cyclic2d;
-48IndexTaskMap task_2 linearize3D;
-49IndexTaskMap task_2 linearize2D;
-50IndexTaskMap task_3 block1d;
-51IndexTaskMap task_3 cyclic1d;
-52IndexTaskMap task_3 cyclic2d;
-53IndexTaskMap task_3 linearize3D;
-54IndexTaskMap task_3 linearize2D;
-55IndexTaskMap task_5 block1d;
-56IndexTaskMap task_5 cyclic1d;
-57IndexTaskMap task_5 cyclic2d;
-58IndexTaskMap task_5 linearize3D;
-59IndexTaskMap task_5 linearize2D;
-60
-61m_2d = Machine(GPU);
-62def same_point(Task task) {
+19}
+20
+21IndexTaskMap task_1 linearize2D;
+22IndexTaskMap task_2 linearize2D;
+23IndexTaskMap task_3 linearize2D;
+24IndexTaskMap task_5 linearize2D;
+25
+26m_2d = Machine(GPU);
+27def same_point(Task task) {
 28 return m_2d[*task.parent.processor(m_2d)];
-64}
+29}
 Figure A7: For Solomonik’s algorithm, we show the mapper produced by the mapper agent at iteration 10.